### v1

# Decoding Quantum LDPC Codes using Collaborative Check Node Removal

Abstract
The fault tolerance of quantum devices requires on-par contributions from error-correcting codes and suitable decoders. One of the most explored error-correcting codes is the family of Quantum Low-Density Parity Check (QLDPC) codes. Although faster than many of the reported decoders for QLDPC codes, iterative decoders fail due to the colossal degeneracy and short cycles intrinsic to these codes. We present a strategy to improve the performance of the iterative decoders based on a collaborative way to use the message passing of the iterative decoders and check node removal from the code’s Tanner graph. We use the concept of bit separation and generalize it to qubit separation. This gives us a metric to analyze and improve the decoder’s performance towards harmful configurations of QLDPC codes. We present a simple decoding architecture to overcome iterative decoding failures by increasing the separation of trapped qubits without incurring any significant overhead.
I Introduction
Quantum Low-Density Parity Check (QLDPC) codes stand as the most suitable candidate for achieving practical fault tolerance with their low overhead [1] and good error correction properties [2, 3]. A good error-correcting code refers to code that has a constant rate and shows linear or constant distance scaling. In comparison with the classical LDPC codes, QLDPC code constructions are less trivial. The first set of construction towards acheiving the same was the hypergraph product (HGP) codes by Tillich and Zémor [3]. These QLDPC codes achieve a constant rate and a quadratic distance scaling (that is, $d\propto\sqrt{n}$) as an outcome of the tensor product between any two classical codes. The well known Toric code is a class of HGP code itself. More recent studies have presented the first constructions of asymptotically good QLDPC codes [4, 5]. For instance, Panteleev and Kalachev first proposed a generalization of the HGP construction [6], which later they renamed to Lifted Product codes [2], reporting the first QLDPC codes with almost linear distance.
The performance of these QLDPC codes are far from optimal under the current decoders available. Unlike the classical LDPC codes, high degeneracy and short cycles present in the Tanner graph of the code poses a big hurdle during the decoding. Particularly iterative Belief Propagation (BP) based decoders completely fail while decoding these highly degenerate quantum codes. The state-of-the-art QLDPC decoder with a mammoth amount of success rate is the BP-OSD decoder[6]. It uses a post processing classical method based on the Ordered Statistics Decoding (OSD), to solve an inverse based problem to exhaustively find minimum weight error patterns that satisfy the syndrome. Although very accurate, it has a $\mathcal{O}(n^{3})$ time complexity in the number of physical qubits $n$ used by the error correcting code. Although, more accurate and higher order decoding of the same is too expensive for using in any practical cases. Recent attempts are made to achieve a similar accuracy with reduced time complexity [7, 8]. Also there have been certain decoders which in part, use BP and leverages from the properties of large codes, such as expansion to post-process the trapped BP outputs [9].
Different harmful configurations intrinsic to the QLDPC codes are considered as the main sorce of iterative decoding failures. Previous attempts were made based on the knowledge of these configurations and helping the decoder learn or facilitate asymmetric message passing rules to avoid failures [10, 11, 12]. In this work, we present a simple approach to assist the iterative decoder without the knowledge of these configurations. In this way we save a lot of effort towards exhaustive analysis and expensive learning based resource exploitation. We provide a notion of the qubit separation and show that removing certain mis-satisfied and symmetric check nodes (i.e. stabilizers) can improve the iterative decoder’s performance. We show that the decoder’s performance is directly linked with increased separation of the qubits belonging to the harmful configuration sets. We also leverage from the intrinsic degeneracies to estimate the separation of qubits belonging to symmetric stabilizers sets. We show a probabilistic check node removal approach like [13] offers highly separated qubits and thus can ensure a better performant decoding architecture than that of the decoders involving only iterative message passing. We observe that the proposed collaborative architecture incurs very little overhead to the overall decoding process compared to the existing post-processing methods and offers a performance bridge between iterative and the iterative + post-processing-based decoders.
The paper is organized as follows: In Section II, we describe the classical and quantum error correction process and also provide necessary details of the Generalized Hypergraph Product (GHP) code that we use primarily for testing and analyzing the decoder. In Section III, we analyze the harmful configurations that arise for GHP quantum code resulting in decoding failures and discuss how one can break pass those failures by increasing the separation of the trapped qubits. Finally, we discuss the proposed collaborative decoding in Section IV and show that the proposed sub-decoding mode can indeed improve from the trapped error floor region of a BP min-sum decoder and compare the performance of the proposed work with BP and BP-OSD0 using some memory experiments for the $[[882,24]]$ and $[[1270,28]]$ GHP codes.
II Preliminaries
We now briefly discuss the classical and quantum error correction process and summarize the min-sum BP decoder.
II-A Classical Error Correction
Classical error correcting codes are a typical binary map, which encodes $k$ bits of information into $n$ bits of codeword $\mathrm{c}$, where $n>k$. The error correction mechanism can be expressed through the generator matrix $\mathbf{G}$ or the parity check matrix $\mathbf{H}$. The former method uses the fact that the rows of $\mathbf{G}$ span the set of all codewords $\mathcal{C}$. The latter defines a code through the null space of $\mathbf{H}$, that is, any codeword $\mathrm{c}\in\mathcal{C}$:
$\mathbf{H}\mathrm{c}^{T}=0\,\,\text{mod}\,\,2.$ (1)
This establishes the foundation to identify and correct errors. For example, if there is an error $\eta$ occurred, the codeword becomes erroneous, i.e., $\mathrm{c}\rightarrow\mathrm{y}:=\mathrm{c}+\eta$ and therefore Eq. (1) gets violated. Knowledge of the error $\eta$ is often deciphered using the syndrome $\mathrm{s}$ computed using $\mathrm{y}$ and $\mathbf{H}$ as
$\mathbf{H}\mathrm{y}^{T}=\mathrm{s}.$ (2)
There exist nonzero syndromes only for errors $\eta$ with a Hamming weight upper bounded by $\text{wt}(\eta)<d$, where $d$ is the minimum distance of the classical code. A decoder typically leverages the syndrome information and predicts the error occurred but the performance of the decoder is often restricted by both the decoding algorithm as well as the constraints imposed by the minimum distance of the code.
II-B Quantum Error Correction
Quantum error correcting codes similar to classical codes use redundancy to encode the information of $k$ logical qubits into $n$ physical qubits. The fundamental errors that qubits suffer include continuous errors, unlike the bit flips in the classical case. A typical error in a quantum system can be expressed as a linear combination of errors in the Pauli set $\{I,X,Y,Z\}$. The Pauli-$X$ operator acts similarly to the classical bit flip error by flipping the bit-state of a qubit. Meanwhile, the Pauli-$Z$ operator changes the phase of a quantum state. The encoded logical qubit state |$\ket{\psi}_{L}$ belongs to the space of $(+1)$ eigenstates of a set of Pauli operators $\mathcal{S}\subset\mathcal{G}_{n}$, called the stabilizers of the code. Any error must anti-commute with one of the stabilizers for the decoder to be able to detect it. A one-to-one map exists between the stabilizer set and parity checks matrices for a quantum code where the rows map to the stabilizers of the code. For instance, if a quantum $[[n,k,d]]$ code has a stabilizer set $\mathcal{S}=\{s_{1},s_{2},...,s_{n-k}\}$. The parity check matrix can be defined by stacking the stabilizers in the rows to form the matrix
$\begin{bmatrix}s_{1}\\ s_{2}\\ \vdots\\ s_{n-k}\end{bmatrix}.$
The symplectic format allows a binary representation of the parity check matrix of the above representation. In this format, each of the $n$ qubit Pauli operators is represented using a $2n$ bit string. For example, one qubit Pauli operator follows
$X\mapsto(1|0),$ (3)
$Y\mapsto(1|1),$ (4)
$Z\mapsto(0|1).$ (5)
The parity check matrix can be described in binary form as:
$\mathbf{H}=\begin{bmatrix}\mathbf{H}_{X}|\mathbf{H}_{Z}\end{bmatrix}.$ (6)
For any valid code, the stabilizers must all commute with each other, requiring the following to hold:
$\mathbf{H}_{X}\mathbf{H}^{T}_{Z}+\mathbf{H}_{Z}\mathbf{H}^{T}_{X}=0.$ (7)
This constrains the development of quantum codes. Particularly in this work, we focus on a very well-studied class of quantum codes called quantum low-density parity check (QLDPC) codes [14]. More specifically, we focus on the class of Calderbank-Shor-Steane (CSS) codes [15, 16] with a special property that the number of $1$’s in each row and column is upper bounded by a constant with the number of qubits going to infinity. This results in a very sparse parity check matrix, which is very efficient with respect to iterative decoding. The stabilizers of these codes are of the independent type, meaning they contain either Pauli-$X$ or Pauli-$Z$ operators and as a result, Eq. (6) takes a special form
$\mathbf{H}=\begin{bmatrix}\mathbf{H}_{X}&\mathbf{0}\\ \mathbf{0}&\mathbf{H}_{Z}\end{bmatrix}.$ (8)
This ensures that decoding can be performed independently for syndromes due to bit flips and the syndromes due to phase flips. Also, Eq. (7) can now be reduced to
$\mathbf{H}_{X}\mathbf{H}^{T}_{Z}=0.$ (9)
Following this, we explore a special generalization of the HGP codes, with improved code properties that require specialized analysis.
II-C Generalized Hypergraph Product (GHP) Codes
An intuitive way to construct QLDPC codes based on Eq. (9), is using two commuting matrices and stack them together to form the parity check matrix of the QLDPC code [17]. Panteleev and Kalachev use this intuition to propose a generalization of the standard hypergraphproduct codes [6]. In their work they consider two square matrices, elements of which belongs to a ring called ring of circulants defined below in II.1.
Definition II.1 (Ring of Circulants).
A ring of circulant $\mathbb{A}_{L}$ satisfies all the properties of a ring with its members being the circulant permutation matrices or sometimes referred to as the polynomials. The whole ring is characterized by a parameter called "lift", which is essentially the dimension of the permutation matrices.
If the ring of circulants contains only null and identity, then the GHP CSS code reduces to HGP code. The parity check matrix of a GHP code is of the following form:
$\mathbf{H}_{X}=[A,bI_{m}]\,\,\text{and}\,\,\mathbf{H}_{Z}=[b^{T}I_{n},A^{T}],$ (10)
where A is a matrix $A=[a_{ij}]_{m\times n}$, such that $a_{ij}\in\mathbb{A}_{L}$ and $m=n$. Further, $b$ is a polynomial in $\mathbb{A}_{L}$ and $I_{m}$ and $I_{n}$ are diagonal matrices with its diagonal entries being the identity of the ring. For example, we will extensively use the $[[882,24]]$ code discussed in [6], which uses
$A$ $=\begin{pmatrix}x^{27}&0&0&0&0&1&x^{54}\\ x^{54}&x^{27}&0&0&0&0&1\\ 1&x^{54}&x^{27}&0&0&0&0\\ 0&1&x^{54}&x^{27}&0&0&0\\ 0&0&1&x^{54}&x^{27}&0&0\\ 0&0&0&1&x^{54}&x^{27}&0\\ 0&0&0&0&1&x^{54}&x^{27}\end{pmatrix},$ (11)
$b$ $=\begin{pmatrix}1+x+x^{6}\end{pmatrix}I_{7},$ (12)
where $x^{i},\,\,\forall i\in\{1,...,L-1\}$ are the polynomials with it’s columns shifted by $i$ from the identity. Also the lift for the ring is set as $L=63$, thus all this polynomials are $(63\times 63)$ matrices over $\mathbb{F}_{2}$.
II-D Iterative Decoding
We discussed in the previous sections that errors anti-commuting with the stabilizers result in non-zero syndrome and indicate the occurrence of errors. Based on the received syndrome $\mathrm{s}$, the decoding problem is a search for the error, which maximizes the posterior probability:
$\hat{e}=\operatorname*{arg\,max}_{e\in\mathcal{G}_{n}}P(e|\mathrm{s}),$ (13)
where, $\hat{e}$ is the best estimate of the error $e$ and $\mathcal{G}_{n}$ is the set of all $n$-qubit Pauli operators. In case of quantum error correction, the optimal decoding is instead solving the following optimization:
$\hat{e}=\mathrm{any}(E):E=\operatorname*{arg\,max}_{E\in\mathcal{E}}P(E|\mathrm{s}),$ (14)
where $E$ is the coset of an error $e\in\mathcal{G}_{n}$ such that $E:=e\mathcal{S}$. $\mathcal{E}$ is a collection of all such possible cosets. The estimated error $\hat{e}$ of Eq. (14) can be any operator from the maximum probability coset $E$, as members of a single coset has the same logical effect on the state.
Throughout this article we use iterative decoding and BP decoding in an equivalent manner. BP is an iterative message-passing algorithm due to its continuous back-and-forth information-passing characteristics over a bipartite Tanner graph of the code. BP primarily evaluates the marginals of Eq. (13), i.e.
$\hat{e}_{i}=\operatorname*{arg\,max}_{e_{i}\in\{e_{1},e_{2},....,e_{n}\}/e_{i}}P(e_{i}|\mathrm{s}),$ (15)
Therefore it so happens that the maximized $\hat{e}$ in (13) does not always satisfy Eq. (14). Many variants of the BP algorithm exist, primarily depending on the application of the algorithm. We describe very briefly the min-sum algorithm, which we used for the numerical experiments and also offers better numerical stability than the others. This algorithm performs series of binary message passing from data qubit nodes to it’s adjacent check nodes and vice-versa, producing soft outputs for each data qubit nodes. We use $u_{i}$ and $v_{j}$ to denote the $i^{th}$ data and $j^{th}$ check nodes respectively. Also, we assume the notation $m^{k}_{a\rightarrow b}$ to portray message passed at the $k^{th}$ iteration of the algorithm from node $a$ to node $b$ on the Tanner graph. Following we describe an overall brief summary of the standard message passing rule of the min-sum BP:
  1. 1.
    Initialization: Initially from each data qubits a message $m^{0}_{u_{i}\rightarrow v_{j}}$ is passed as follows:
    $m^{0}_{u_{i}\rightarrow v_{j}}=\lambda_{i}=\log(\frac{1-p_{i}}{p_{i}});\,\,\forall i,j:\mathbf{H}_{ji}=1.$ (16)
  2. 2.
    Check to data qubit messages: For the subsequent iterations, message $m^{k}_{v_{j}\rightarrow u_{i}}$ is passed from each check node to data qubits. Although note we exclude the message passing to the qubit node, which was immediately involved in message passing at ${k-1}^{th}$ iteration. We define a quantity $w=\min_{u_{p}:\,N(v_{j})/u_{i}}\{|m^{k-1}_{u_{p}\rightarrow v_{j}}|\}$ and the message passed is as follows:
    $m^{k}_{v_{j}\rightarrow u_{i}}=(-1)^{\mathrm{s}_{j}}\alpha\left(\prod_{u_{p}:\,N(v_{j})\backslash u_{i}}sign(m^{k-1}_{u_{p}\rightarrow v_{j}})\right)w,$ (17)
    where, $\alpha$ is a scaling factor and can affect the convergence and $\mathrm{s}_{j}$ is the syndrome value corresponding to the check node $v_{j}$.
  3. 3.
    Data qubit to check messages: Except the initialization as described, at each iteration $k$, the messages passed from data qubits to check nodes follows:
    $m^{k}_{u_{i}\rightarrow v_{j}}=\lambda_{i}+\sum_{v_{p}:N(u_{i})\backslash v_{j}}m^{k-1}_{v_{p}\rightarrow u_{i}}$ (18)
  4. 4.
    Hard decision: At the end of each iteration, a soft output is evaluated for each qubit node. The soft output represents the collective response received from all the check nodes and is as follows:
    $\gamma_{i}\leftarrow\lambda_{i}+\sum_{v_{p}:N(u_{i})}m^{k}_{v_{p}\rightarrow u_{i}}.$ (19)
    If $\mathrm{sign}(\gamma_{i})=-1$, then the hard decision for data qubit $i$ is $\hat{e}_{i}=1$, otherwise $\hat{e}_{i}=0$.
The whole process continues until all the unsatisfied checks are satisfied or a pre-determined number of message-passing iterations is reached, and the syndrome is still non-zero. In the former, the decoding is said to have succeeded, and the latter reflects a decoding failure. A cycle-free Tanner graph guarantees the success of BP within a finite number of iterations, although QEC offers plenty of short cycles along with inevitable degeneracies, which makes BP fail miserably. In the following section, we discuss one such case where some inherent configurations of the GHP code render BP an ineffective decoder.
$v_{0}$$c_{0}$$c_{1}$$v_{1}$$c_{2}$$c_{7}$$c_{6}$$v_{6}$$c_{12}$
(a) Classical-type Trapping Set
$v_{0}$$c_{0}$$v_{477}$$c_{405}$$v_{405}$$c_{406}$$v_{478}$$c_{352}$$v_{351}$$c_{357}$$v_{483}$$c_{6}$$c_{1}$$c_{411}$$c_{351}$
(b) Quantum Trapping set
Figure 1: In (a) we show a typical $(3,3)$ 6-cycle classical type trapping set and in (b) is a $(6,0)$ Quantum Trapping set of the $[[882,24]]$ GHP code.
III Trapping Sets and Separation Vectors of GHP codes
Iterative message passing-based decoders are vulnerable when it comes to certain error configurations. These are heavily studied and formally named as the Trapping sets (TS) and we will use $\mathcal{T}_{s}$ to denote the set of data qubits that are trapped. If $\mathcal{T}_{s}$ contains $a$ data qubits and $b$ odd degree stabilizer checks then it is a $(a,b)$ trapping set of the code. We use the notation $V_{1}$ to describe the set of qubits $v\in\mathcal{T}_{s}$, such that $v$ is adjacent to odd degree stabilizer checks. Note that there can be many possible configurations for any $(a,b)$ TS, depending on the geometry of the code and the decoder under consideration.
III-A Trapping set analysis
For QLDPC codes there exists two class of harmful TS configurations: Classical- type trapping sets (CTS) and Quantum trapping sets (QTS). The definitions of CTS and QTS are briefly covered in III.1 and III.2, which we adopt from [11].
Definition III.1 (Classical-type Trapping sets).
A classical-type trapping set is a set of qubits which either do not converge or they are adjacent to unsatisfied stabilizer checks after a predetermined number of iterations of a syndrome based iterative decoder.
For example in Fig. 1(a), we show a $(3,3)$ CTS obtained from the $\mathbf{H}_{Z}$ of the same $[[882,24]]$ GHP code mentioned in (11). As can be seen from the qubit and check node indexing, the CTS is formed only due to the contributions of a particular circulant matrix, namely $b^{T}I_{n}$. This type of CTS formation is also discussed in [18]. The min-sum iterative decoder in II-D fails to converge in this scenario due to the symmetric message passing rules. For example, for a received syndrome $\mathrm{s}=\{c_{0},c_{1},c_{2},c_{6},c_{7},c_{12}\}$, the min-sum BP based decoder oscillates between an all-zero error pattern and an error pattern of $\{v_{0},v_{1},v_{6}\}$, as shown in Fig 2.
$v_{0}$$c_{0}$$c_{1}$$v_{1}$$c_{2}$$c_{7}$$c_{6}$$v_{6}$$c_{12}$100010001
(a) Iteration $\mathcal{K}=1$
$v_{0}$$c_{0}$$c_{1}$$v_{1}$$c_{2}$$c_{7}$$c_{6}$$v_{6}$$c_{12}$000000000
(b) Iteration $\mathcal{K}=2$
$v_{0}$$c_{0}$$c_{1}$$v_{1}$$c_{2}$$c_{7}$$c_{6}$$v_{6}$$c_{12}$100010001
(c) Iteration $\mathcal{K}=3$
Figure 2: We show a typical oscillatory behaviour of the min-sum BP algorithm (II-D) for the $(3,3)$ CTS of the $[[882,24]]$ GHP code. The square colour nodes represent the input unsatisfied check nodes. The final hard information passed to each of the data qubits $v_{j}\in V_{1}$ is shown for each complete iteration of the algorithm. Qubits $v_{j}\notin V_{1}$ are not affected by the message passing and thus not highlighted. The algorithm is initialized with errors indicated at $v_{0}$, $v_{1}$ and $v_{6}$. Thereafter, a non-convergent oscillatory behaviour between all zero and $\{v_{0},v_{1},v_{6}\}$ is observed.
The other type of TS that often occurs in the QLDPC codes are the Quantum Trapping sets (QTS) as defined in III.2. These trapping sets do not have any odd-degree check nodes. Thus, a QTS with $a$ number of data qubits is referred as a $(a,0)$ trapping set. Fig. 1(b) shows a typical QTS formed for the same GHP code with $6$ data qubits in $\mathcal{T}_{s}$.
Definition III.2 (Quantum Trapping sets).
A quantum trapping set is a collection of an even number of qubits and symmetric stabilizers, such that the induced sub-graph contains no odd-degree stabilizer check nodes. This type of trapping set can always be partitioned into isomorphic disjoint subsets with a set of common odd-degree stabilizer (check) neighborhoods.
The two disjoint symmetric subsets of the $(6,0)$ TS, can be identified as $V_{a}=\{v_{0},v_{351},v_{405}\}$ and $V_{b}=\{v_{477},v_{478},v_{483}\}$. Further, as per the definition III.2. $V_{a}$ and $V_{b}$ have the same set of odd-degree stabilizer check neighborhood. That is, $N(V_{a})=N(V_{b})=\{c_{0},c_{1},c_{6},c_{405},c_{406},c_{411},c_{351},c_{352},c_{367}\}$. Following the Lemma 1 from [11], error patterns of cardinality less or equals $3$ on any of the subset is a harmful configuration for the $(6,0)$ TS. For instance any $e_{a}\subseteq V_{a}:|e_{a}|\leq|V_{a}|$ has an exact twin $e_{b}\subseteq V_{b}$. This results into an oscillation between $e_{a}\oplus e_{b}$ and an all zero error at each iterations of the min-sum decoder discussed in II-D.
Lemma 1.
Iterative decoders with critical number $\frac{a}{2}$ posses no TS characteristics for the $(a,0)$ QTS, if the cardinality of the error subsets $e_{a}\subseteq V_{a}$ or $e_{b}\subseteq V_{b}$ exceeds $\frac{a}{2}$.
Note that, unlike CTS, the common stabilizer neighborhood of the $(6,0)$ QTS is itself the set of odd degree check nodes for each of the disjoint subsets and therefore emerges as the unsatisfied check nodes satisfying $\mathtt{UNSAT}\subseteq N(V_{a})=N(V_{b})$. This requires different measures to approach how the iterative message passing can be improved.
III-B Use of qubit “separation" to improve iterative decoders
We now show that the iterative decoder can leverage the increased “separation” of the trapped data qubits. We observe that increasing the separation of data qubits has a direct implication towards the decoder’s performance. The oscillatory behavior of the TS discussed in the previous section is majorly contributed by the mis-satisfied check nodes of the TS. For example the $(3,3)$ TS in Fig 1(a)- $c_{1}$, $c_{6}$ are the mis-satisfied check nodes for trapped $v_{0}$. The term mis-satisfied justifies the dilemma they produce as they although are satisfied, passes mis-information, resulting in a non-convergence scenario for $v_{0}$. This phenomenon is also reflected in Fig 2. This leads to the intuition behind the separation of a qubit. It is essentially a measure of how far an incorrect trapped qubit is from another trapped qubit in the computation tree [19]. The core notion behind this measure is to identify the mis-satisfied check nodes vulnerable to decoding.
The computation tree (CT) is often used to analyze the performance of iterative decoders in classical settings [20, 19]. Consider a trapping set $\mathcal{T}_{s}$, with a set of odd degree data qubits $V_{1}$. It is built either from a trapped qubit $v\in V_{1}$ or from an unsatisfied check node $c\in C_{1}$ as a root. From the root, we add descendant check and data qubits according to the iterative message passing (or, the Tanner graph of the code). We denote a CT for $\mathcal{K}$ iterations of complete message passing cycles (i.e., qubit $\rightarrow$ check $\rightarrow$ qubit), with a qubit $v$ as root by $T_{\mathcal{K}}(v)$, and if the root is an unsatisfied check node $c$, then it is denoted by $T_{\mathcal{K}}(c)$. Therefore, $\mathcal{K}$ also indicates the level of the computation tree and from here onward we will alternately use it to denote iteration of the min-sum decoder as well as the level of a computation tree. Based on the computation tree of the trapped qubits for both CTS and QTS, the separation of the same is calculated, described in definition III.3 and III.4.
Definition III.3 (Qubit separation for CTS).
If an erroneous data qubit node $v\in V_{1}$ for any $V_{1}\subset\mathcal{T}_{s}$; has atleast one degree-one check node $c$, such that, within $\mathcal{K}$ number of message passing iterations of the decoder, there exists no more data qubit $u\in V_{1}$ as a descendant of $c$ in $T_{\mathcal{K}}(c)$, then the root qubit $v$ is said to be $\mathcal{K}$ separated.
In Fig. 5 of the appendix, we show the computation tree $T_{4}(v_{0})$ for the $(3,3)$ CTS. We can see that the qubit $v_{0}$ has a separation of $\mathcal{K}=2$, as at $\mathcal{K}=3$ there exists $v_{6}\in V_{1}$. Also, if there are $a$ qubits in $\mathcal{T}_{s}$, then separation of each of those qubits represented in an one dimensional vector of length $a$ is called the separation vector of that $\mathcal{T}_{s}$. Note at level $\mathcal{K}=3$ of $T_{4}(v_{0})$, if we remove $c_{6}$, the separation of $v_{0}$ is increased to $3$, i.e. $\mathcal{K}=3$ for $v_{0}$. Thus, an increase in separation is intrinsic to the removal of mis-satisfied check nodes. This in-turn helps provide good information to the iterative decoders, see appendix VI-A for details. For QTS, we propose a slightly different definition for the qubit separation.
Definition III.4 (Qubit separation for QTS).
If a QTS $\mathcal{T}_{s}$ has degenerate subsets $V_{a}$ and $V_{b}$, then for an erroneous data qubit node $v\in V_{a}$ with atleast one degree-one check node $c$, is said to be $\mathcal{K}$ separated, if within $\mathcal{K}$ number of message passing iterations of the decoder, there exists no more data qubit $u\in V_{a}$ as a descendant of $c$ in $T_{\mathcal{K}}(c)$.
Similarly, we define the separation for the qubits in $V_{b}$. We note that for QTS, the contributions of the harmful information are channelled through the check nodes connecting the disjoint symmetric subsets. These check nodes play the same role as the mis-satisfied check nodes of the CTS. Although they themselves can be the only unsatisfied check nodes for the harmful error configurations, as we discussed before. In Appendix VI-A, we show for $(6,0)$ QTS, removal of check node pairs from $T_{k}(v_{i})$ (in accordance with definition III.4) has an impact in increasing the separation for $v_{i}$, where $v_{i},v_{j}\in\{V_{a},V_{b}\}$. Further, we provide evidences in appendix VI-A, showing that removal of these checks help the iterative decoder predict correct syndromes, which was not possible through the iterative decoder under normal consitions.
In the following section, we first draw an analogy with the classical results from Zhang et al. [21] to show that increased separation of a trapped node successfully breaks typical unsuccessful decoding scenarios of the iterative decoder. Further, we perform numerical memory experiments to:
  1. 1.
    show the advantage of removing mis-satisfied check nodes through a collaborative decoding approach,
  2. 2.
    benchmark the same against min-sum BP and state-of-the-art BP+OSD decoder.
IV Decoding via Collaborative Check Node Removal
As discussed in the previous section, removing the correct check node from the computation tree of a trap node increases its separation and improves the performance of iterative decoders. However, knowing the exact level in the CT from where the check node removal can be done requires the TS information. For example, Fig. 5 in Appendix VI-A shows clearly that the candidate check nodes for removal are in level $\mathcal{K}=3$ of the CT $T_{4}(v_{0})$. These candidate check nodes are the same mis-satisfied check nodes, information of which can only be known from the TS. Thus, a natural practical way to approach this scenario is through the same probabilistic approach of [13], which increases the separation of the trapped nodes without knowing the particulars of the TS of the quantum code. From here onward, we denote neighbor of a qubit and check as $N(v)$, $N(c)$ respectively and alternately use $t$ and $\mathcal{K}$ for algorithmic syntax purposes.
IV-A Weak check node removal
From here on we address the decoding algorithm for the CTS. Later, we show generalization to QTS requires modification in the removal rate of check nodes. We saw before that the CTS for the GHP code is rooted in the contribution of a single circulant matrix. For example, the CTS shown in Fig 1(a) is due to the circulant matrix $bI_{7}$. These types of trapping sets have odd-degree check nodes adjacent to them, which we exploit to initialize the decoding. Therefore, we identify the decoding problem as syndrome based and extend the definition of CT construction with unsatisfied check nodes as the roots. Further, we adopt the notion of week check nodes from [13], which assumes that at any level of $T_{\mathcal{K}}(c),\forall c\in\mathtt{UNSAT}$; there exist at least one weak check node, which is defined in IV.1.
Definition IV.1 (Weak check nodes).
If in level $t$ of the computation tree $T_{\mathcal{K}}(c)$ has a root unsatisfied check node $c\in UNSAT$, then there exists at-least a check node $z$, which has the shortest path to any of the nearest qubit $v\in\mathcal{T}_{s}$, then $z$ is a weak check node at level $\mathcal{K}=t$ of $T_{\mathcal{K}}(c)$.
Therefore, instead of removing the exact mis-satisfied check nodes from $T_{\mathcal{K}}(v)$, we remove the weak check nodes from level $\mathcal{K}=1$ of $T_{\mathcal{K}}(c_{i})$. It also improves the separation of the adjacent trapped qubit of the unsatisfied check node $c_{i}$. For example, Fig 5 in Appendix VI-A shows $c_{5}$ as the weak check node at $\mathcal{K}=1$. Even if the knowledge of $(3,3)$ CTS is unknown, simply removing $c_{5}$ from $\mathcal{K}=1$ of $T_{k}(c_{0})$ can increase the separation of the adjacent qubit $v_{0}$. This helps introduce a lot of freedom and simplicity to the decoding.
Obviously one can remove weak check nodes from any level $\mathcal{K}$ of $T_{\mathcal{K}}(c_{i})\,\,\forall c_{i}\in\mathtt{UNSAT}$. Also it is obvious that not all the check nodes at level $t$ are the weak check nodes and removing these will have a negative affect in the further iterative decoding. Therefore, we use the parameter “deselection degree” denoted as $\mathtt{df}$, to restrict the number of random check node removal from any level of the CT. Also, note that alternatively one can identify the level $\mathcal{K}=t$ of the computation tree as the level of all leaf check nodes of the computation tree $T_{\mathcal{K}}(c_{i})$. This method of removing weak leaf nodes randomly is the key assistive part of the overall decoding architecture. The complete algorithm is known as the Fixed Node Check Removal (FNCR) [13], which is described in Algorithm 1.
input : $\mathbf{H}$, $\mathtt{UNSAT}$, $\mathtt{t}$, $\mathtt{df}$
output : Modified parity check matrix $\mathbf{H}_{FN}$
$\mathbf{H}_{FN}=\mathbf{H}$;
  // Initialization
1 $\mathtt{rem}$ = [];
2for $\mathtt{check}$ in $\mathtt{UNSAT}$ do
3       $\mathtt{leaf}$leftarrow All the leaf nodes of $T_{t}(\mathtt{check})$;
4       $\mathtt{dis}$leftarrow random.chose($\mathtt{leaf}$,$\mathtt{df}$);
5       $\mathtt{rem}$circ $\mathtt{dis}$;
6foreach $\mathtt{uc}$in $\mathtt{rem}$ do
       $\mathbf{H}_{FN}\leftarrow\mathbf{H}_{FN}\backslash\mathbf{H}[\mathtt{uc}]$;
        // Row removal
7
return $\mathbf{H_{FN}}$ ;
  // The modified matrix
Algorithm 1 FNCR decoding
IV-B Collaborative architecture
For CTS, it is enough to set $\mathtt{df}=1$, as can be seen from Fig. 5 and the numerical results of appendix VI-A. For improving the qubit separation in QTS, ideally, the removal rate should be set to at least $\mathtt{df}=2$, as seen from Fig. 6 in appendix. The modified check matrix output of algorithm 1 has a high probability of having qubits with increased separation and thus offers better convergence with iterative decoders. Therefore, an iterative decoder can be used twice, first with the full parity check matrix and again on the modified parity check matrix from Algorithm 1. We call this approach a collaborative decoding, details of which are described in Algorithm 2. Collaborative decoding can be considered primarily as using a BP decoder in two modes. During the main decoding mode, BP uses the complete parity check matrix of the QLDPC code and the other mode is where it uses the modified parity check matrix output of Algorithm 1, for predicting an error estimate. The later is considered as the sub-decoding mode.
input :
  • Syndrome:= $\mathtt{s}$ and corresponding $\mathbf{H}$,
  • Channel llr values:= $\mathtt{llr}$,
  • Maximum iterations of primary BP:= $\mathtt{max_{iter}}$,
  • Maximum iteraions of sub BP:= $\mathtt{max_{sub}}$,
  • Deselection degree: $\mathtt{df}$,
  • Sampling degree: $\mathtt{sf}$,
  • Maximum FNCR rounds: $\mathtt{fr}$,
  • Tolerance of Main BP iterations: $\mathtt{tol}$.
output : Total predicted error pattern $\mathtt{\hat{e}}$.
$\mathtt{unchanged=itn=0}$;
  // Main decoder init
1 $\mathtt{s_{prev}:=s}$;
[…]
3       $\mathtt{\hat{e}_{bp}}$ = $\mathsf{BP}$($\mathbf{H},\mathtt{llr},\mathtt{max_{iter}}$).$\mathsf{decode}$($\mathtt{s}$);
4       $\mathtt{\hat{s}_{bp}}$ = $\mathtt{\hat{e}_{bp}}\mathbf{H}^{T}$;
5       $\mathtt{itn}\leftarrow\mathtt{itn}+\mathtt{max_{iter}}$;
6       if $\mathtt{\hat{s}_{bp}}=\mathtt{s_{prev}}$ then
7            $\mathtt{unchanged\leftarrow unchanged+1}$
8      else
9            $\mathtt{unchanged=0}$
10      $\mathtt{s_{prev}\leftarrow\hat{s}_{bp}}$;
11
[…]
13 $\hat{e}_{sub}:=\mathsf{zeros}\mathtt{((}\mathtt{fr},\mathbf{H}.\mathsf{shape}\mathtt{[1]}\mathtt{))}$ ;
14 while $\mathtt{i<fr}$ do
15       $\mathtt{UNSAT=[]}$;
16       $\mathtt{s_{fn}\leftarrow s+\hat{s}_{bp}+(\sum_{j}s_{sub}[j]),\forall j<i}$;
17       if $\mathtt{check=0,\forall check\in s_{fn}}$ then
18             return $\mathtt{\hat{e}\leftarrow\hat{e}_{bp}+\sum_{i}\hat{e}_{sub}[i]}$;
19             End;
20
21      for $\mathtt{check}$ $\mathsf{in}$ $\mathtt{s_{fn}}$ do
22             if $\mathtt{check\neq 0}$ then
23                   $\mathtt{UNSAT}$.$\mathsf{extend}$($\mathtt{check}$);
24
25
26      $\mathtt{u_{s}=}\mathsf{random.sample}\mathtt{(UNSAT,sf)}$;
27       $\mathbf{H}_{FN}=\mathsf{FNCR}(\mathbf{H},\mathtt{u_{s}},\mathtt{t},\mathtt{df})$;
28       $\mathsf{BP_{sub}}=\mathsf{BP}(\mathbf{H}_{FN},\mathtt{llr},\mathtt{max_{sub}})$;
29       $\mathtt{\hat{e}_{sub}[i]=}\mathsf{BP_{sub}.decode}\mathtt{(s_{fn})}$;
30       $\mathtt{s_{sub}[i]\leftarrow\hat{e}_{sub}[i]}\mathbf{H}^{T}$;
31       $\mathtt{i}\leftarrow\mathtt{i}+\mathtt{1}$
return $\mathtt{\hat{e}\leftarrow\hat{e}_{bp}+\sum_{i}\hat{e}_{sub}[i]}$
Algorithm 2 Collaborative BP decoding with FNCR
In Fig 3, we show the advantage of using BP in the sub-decoding mode. We observe, out of numerous sampled instances, a case where the main BP decoder gets stuck while decoding the syndrome of an error pattern generated at an independent bit flip physical error rate of $p=0.03$ for the $[[882,24]]$ GHP code after a few iterations. The stuck scenario represents the fact that the main decoder cannot identify anymore correct stabilizer violations. The tail of the plot shows that by activating the sub-decoder, we can predict the syndrome correctly and thus, eventually all the errors are correctly predicted. Therefore, we confirm that removing weak check nodes using Algorithm 1 increases the efficacy of the decoding output.
Figure 3: Numerical experiment showing the sub-decoder is able to identify new correct unsatisfied stabilizers, which the main decoder alone can not predict. We observe the decoding performance for the syndrome obtained at independent bit flip physical error rate $p=0.03$ for $[[882,24]]$ GHP code. The main BP decoder can not predict any more correct stabilizer violations and gets stuck at $40^{th}$ iteration. Subsequently, through sub-decoding, we call Algorithm 1 repeatedly and BP is applied on the output of the same. We observe repeated calls of Algorithm 1 leads to an improvement to the performance of BP in the sub-decoding mode. We also observe a tail in the plot signifying newly accurate predictions of the unsatisfied stabilizer checks.
Further, we perform some memory experiments to compare the performance of Algorithm 2 with standard BP min-sum decoder and the state-of-the-art BP+OSD decoder. In Fig 4, we show the results, which indicate that the proposed decoding scheme offers significant improvements in the logical error rates over the BP decoder, indicating a potential breakthrough to circumvent the harmful configurations for the iterative BP. Algorithm 2 give logical error rates better than BP but it lags compared to BP+OSD. While BP+OSD uses an exhaustive search strategy to find the most accurate error patterns, Algorithm 2 is devoid of such high-order post-processing steps. There is still a scope to improve accuracy to identify check nodes. For any improvement, we note that the parameter $\mathtt{df}$ can help improve the decoding success rates.
Figure 4: The logical error rates obtained for the GHP codes from [6], under the independent bit flip Pauli channel. We perform the code capacity LER simulations for both cases. In (a) we benchmark with the $[[882,24]]$ GHP code and in (b) we benchmark the $[[1270,28]]$ GHP code. The LERs are plotted within a $95\%$ confidence intervals, which are indicated through the error-bars. In both cases, LERs are improved using Algorithm 2 over iterative min-sum BP.
IV-C Time Complexity
The collaborative decoding essentially has two phases, one consisting of the action of BP and the other consisting the process of modifying the parity check matrix for the sub-decoder. The cost of the latter mainly comes from the process of constructing the computation tree from the unsatisfied check nodes. Now following the concept of weak check nodes, we discussed that a computation tree of level $\mathcal{K}=1$ is enough to asses the possibility of increasing the separation. For a bi-regular $(d_{v},d_{c})$ QLDPC code, such a tree construction follows from adding the $dc$ adjacent data qubits to the root, which is a $\mathcal{O}(n)$ process. Further the leaf check nodes of level $\mathcal{K}=1$, are added in $\mathcal{O}(m\,d_{c})$ time, where $m$ is the number of stabilizers of the code. As QLDPC codes show extreme sparsity, i.e. $d_{c}<<<m$, Algorithm 1 or the process of generating the modified check matrix is linear, i.e. $\mathcal{O}(n)$ for a single unsatisfied stabilizer or check node. If the support of the syndrome is $\mathrm{supp}(s)=||s||$, then the algorithm has an overall time complexity of approximately $\mathcal{O}(||s||n)$. We also point out that the complexity of Algorithm 1 can even be improved through sampling the unsatisfied check nodes from the syndrome, which essentially reduces the cardinality of the input $\mathtt{UNSAT}$ for Algorithm 1. We use the parameter $\mathtt{sf}$, to determine the sampling rate of such unsatisfied check inputs.
The collaborative decoder can offer an almost linear computational complexity considering the syndrome weight is low (or, low physical error rate of qubits) or, a heavy sampling of the unsatisfied check nodes. Therefore, the time complexity of the overall collaborative decoding has an implicit dependence on the support of the received syndrome.
V Conclusion and Future Work
In this work, we propose a collaborative decoding approach for improving the iterative decoding of QLDPC codes through improving qubit separation, with a special focus to GHP codes. We propose a new way to measure the separations of the qubits trapped inside a symmetric stabilizer set. We used this method to go beyond the analysis of qubit separation of the classical trapping sets. Our decoding architecture is free of any post-processing method and can be considered as a two-mode decoder with its main decoder having $\mathcal{O}(n)$ linear complexity. The sub-decoder’s complexity depends on support of the syndrome (an implicit dependence on the physical error rates of the qubits). We discussed for practical situations the complexity of the sub-decoder can be improved as much as almost linear. We showed that inspite of being a post-processing-free decoding the success of the algorithm is a significant improvement compared to the standard iterative min-sum decoder. In future work, we propose on optimizing the current architecture to restore the lost information due to excessive check node removals thereby approaching better logical failure rates.
Data Availability
The data and scripts used for the numerical experiments are available from the authors on reasonable requests.
Acknowledgements
MB acknowledges ldpc-repository by Joschka Roffe [22], for sourcing BP and BP-OSD [23] decoders used in the memory experiments. AR and MB thank IISER Bhopal for providing the computational resources to conduct the exhaustive simulations. MB acknowledges the doctoral research fellowship from IISER Bhopal.