### v2

# Decoding Quantum LDPC Codes using Collaborative Check Node Removal

Abstract
Fault tolerance of quantum protocols require on-par contributions from error-correcting codes and it’s suitable decoders. One of the most explored error-correcting codes is the family of Quantum Low-Density Parity Check (QLDPC) codes. Although faster than many of the reported decoders for QLDPC codes, iterative decoders fails to produce suitable success rates due to the colossal degeneracy and short cycles intrinsic to these codes. We present a strategy to improve the performance of the iterative decoders based on a collaborative way to use the message passing of the iterative decoders and stabilizer check node removal from the quantum code’s Tanner graph. We particularly introduce a notion of “qubit separation", which gives us a metric to analyze and improve the min-sum Belief Propagation (BP) based iterative decoder’s performance towards harmful configurations of QLDPC codes. We further show that an integration of information measurements (IM) for qubits and it’s adjacent stabilizer checks, can be exploited to extract far better performing results from the collaborative decoding architecture compared to it’s classical predecessor. We analyze the performance of the proposed collaborative decoding architecture, in the context of Generalized Hypergraph Product (GHP) codes. We discuss that the collaborative decoding architecture overcomes iterative decoding failures regarding the harmful trapping set configurations by increasing the separation of trapped qubits without incurring any significant overhead.
Index Terms:
Quantum Error Correction, QLDPC codes, Decoders, Belief Propagation, Trapping sets, Qubit Separation, Generalized Hypergraph Product codes.
I Introduction
Quantum Error correction has been a primary focus of many to realize the quantum protocols with high fidelity. Over the years, Quantum Low-Density Parity Check (QLDPC) codes have emerged as the most suitable candidate for achieving practical fault tolerance due to their low overhead [1] and good error correction properties [2, 3]. A good error-correcting code refers to code that has a constant rate and shows linear or constant distance scaling. In comparison with the classical LDPC codes, good QLDPC code constructions are less trivial. The first set of construction towards achieving the same was the Hypergraph Product (HGP) codes by Tillich and Zémor [3]. These QLDPC codes achieve a constant rate and a quadratic distance scaling (that is, $d\propto\sqrt{n}$) as an outcome of the tensor product between any two classical codes. The well-known Toric code is a class of HGP code itself [4]. More recent studies have presented the first constructions of asymptotically good QLDPC codes [5, 6]. For instance, Panteleev and Kalachev proposed a generalization of the HGP construction [7], which later they renamed to Lifted Product codes [2], reporting the first QLDPC codes with almost linear distance.
The performance of these QLDPC codes is far from optimal under the current available decoders. Unlike the classical LDPC codes, the high degeneracy and short cycles present in the Tanner graph of the code pose a big hurdle during decoding. Particularly iterative Belief Propagation (BP) based decoders completely fail while decoding these highly degenerate quantum codes. For instance, the surface code exhibits no threshold under the BP decoding. In this article, we use both the iterative and BP decoders alternately on an equal footing to represent the same class of message-passing-based decoder. The state-of-the-art QLDPC decoder, which has the capability of improving the decoding success with a mammoth leap is the BP+OSD decoder [7]. This particular decoder uses a post-processing classical method based on the Ordered Statistics Decoding (OSD), to solve a matrix inversion problem and would exhaustively find minimum weight error patterns that satisfy the syndrome. Although very accurate, even zeroth-order OSD post-processing has a $\mathcal{O}(n^{3})$ time complexity in the number of physical qubits $n$ used by the error correcting code. Further improvements in the accuracy require higher-order post-processing of the same and are too expensive to use in any practical cases. Recent attempts have been made to achieve similar accuracy with reduced time complexity [8, 9]. In addition, certain decoders have been developed thatat use BP anleverage properties of large codes, such as expansion to post-process the trapped BP outputs [10].
Different harmful configurations, popularly known as Trapping Sets (TS), intrinsic to the QLDPC codes, are considered the main source of iterative decoding failures. Attempts are made based on the knowledge of these configurations to assist the decoder learn or facilitate asymmetric message passing rules to avoid failures [11, 12, 13, 14]. In this work, we present a simple approach to assist the iterative decoder without any active monitoring requirement for the knowledge of these configurations in decoding. Our motivation is to inhibit a set of information flows that might result in convergence failures for any iteration of the message passing iterative decoder. We introduce the notion of qubit separation and evaluate information measurement (IM) values to identify the high risk stabilizer check nodes for a given set of violated stabilizer check nodes, i.e. syndrome. We propose that these high-risk check nodes having high IM values limit the separation of certain data qubits adjacent to it and show that the iterative decoder is unable to converge for this set of data qubits. Therefore, we conclude that removing these high-risk check nodes (i.e. stabilizers) can improve the iterative decoder’s performance, and doing so, we establish that the decoder’s performance is directly linked to increased separation of the data qubits belonging to the harmful configurations.
A similar intuition was adopted in [15], where the authors removed a set of data qubits in each iteration of the BP decoder as a post-processing step. This essentially deactivates some stabilizer check node in each BP iteration, and they execute such deactivation according to the net soft information of the adjacent data qubits. The key difference between our work and [15] is that we do not modify the codespace, i.e. our decoding sub-routines directly remove stabilizer check nodes without any straight exclusion of information passed from the data qubits to other check nodes. Also, we will see that our proposed decoder turns off multiple stabilizer check nodes in a single iteration, whereas the stabilizer inactivation in [15] is done one by one for each iteration of BP. This, on average, converges the stuck message passing BP quickly, which is further supported by the additional sub-routine of information measurement and quantified through improvement of qubit separation. The proposed way of improving the iterative decoder’s performance saves a lot of effort towards exhaustive analysis and expensive learning based resource exploitation. We also discuss improvements over the probabilistic check node removal approach of Kang et al. [16] and discuss how the proposed algorithmic sub-routines ensure highly separated qubits and thus can offer a better-performing decoding architecture than that of the decoders involving only iterative message passing. We observe that the proposed collaborative architecture incurs very little overhead to the overall decoding process compared to the existing post-processing methods and still offers significant success rate improvements compared to standard iterative decoders like min-sum BP.
The paper is organized as follows: In Section II, we describe the classical and quantum error correction process and also provide necessary details of the Generalized Hypergraph Product (GHP) code that we use primarily for testing and analyzing the decoder. In Section III, we analyze the harmful configurations that arise for GHP quantum code resulting in decoding failures and discuss how one can break pass those failures by increasing the separation of the trapped qubits. Finally, we discuss the proposed collaborative decoding in Section IV and show that the proposed sub-decoding mode can indeed improve from the trapped error floor region of a min-sum BP decoder and compare the performance of the proposed work with BP and BP+OSD00 using memory experiments for the $[[882,24]]$ and $[[1270,28]]$ GHP codes. Additionally, we also provide proof of existing thresholds for the Generalized Bicycle (GB) codes with the use of the proposed decoder.
II Preliminaries
We now briefly discuss the classical and quantum error correction process and summarize the min-sum BP decoder.
II-A Classical Error Correction
Classical error correcting codes are a typical binary map, which encodes $k$ bits of information into $n$ bits of codeword $\mathrm{c}$, where $n>k$. The error correction mechanism can be expressed through the generator matrix $\mathbf{G}$ or the parity check matrix $\mathbf{H}$. The former method uses the fact that the rows of $\mathbf{G}$ span the set of all codewords $\mathcal{C}$. The latter defines a code through the null space of $\mathbf{H}$, that is, any codeword $\mathrm{c}\in\mathcal{C}$ satisfy:
$\mathbf{H}\mathrm{c}^{T}=0\,\,\text{mod}\,\,2.$ (1)
This establishes the foundation to identify and correct errors. For example, if there is an error $\eta$ occurred, the codeword becomes erroneous, i.e., $\mathrm{c}\rightarrow\mathrm{y}:=\mathrm{c}+\eta$ and therefore Eq. (1) gets violated. Knowledge of the error $\eta$ is often deciphered using the syndrome $\mathrm{s}$, calculated using $\mathrm{y}$ and $\mathbf{H}$ as
$\mathbf{H}\mathrm{y}^{T}=\mathrm{s}.$ (2)
There exist non-zero syndromes only for errors $\eta$ with a Hamming weight upper bounded by $\text{wt}(\eta)<d$, where $d$ is the minimum distance of the classical code. A decoder typically leverages the syndrome information and predicts the error that occurred, but the performance of the decoder is often restricted by both the decoding algorithm and the constraints imposed by the minimum distance of the code.
II-B Quantum Error Correction
Quantum error correcting codes, similar to classical codes, use redundancy to encode the information of $k$ logical qubits into $n$ physical qubits. The fundamental errors that qubits suffer include a continuous set of errors, unlike the classical case, where the only source of errors is bit flips at discrete positions. A typical error in a quantum system can be expressed as a linear combination of errors in the Pauli set $\{I,X,Y,Z\}$. The Pauli-$X$ operator acts similarly to the classical bit flip error by flipping the state of a qubit. Meanwhile, the Pauli-$Z$ operator changes the phase of a quantum state. The encoded logical qubit state |$\ket{\psi}_{L}$ belongs to the space of $(+1)$ eigenstates of a set of Pauli operators $\mathcal{S}\subset\mathcal{G}_{n}$, called the stabilizers of the code. Any error must anti-commute with one of the stabilizers for the decoder to be able to detect it. A one-to-one map exists between the stabilizer set and the parity check matrix (PCM) of a quantum code. Each row of the PCM maps to a stabilizer of the code. For instance, if a quantum $[[n,k,d]]$ code has a stabilizer set $\mathcal{S}=\{s_{1},s_{2},...,s_{n-k}\}$. The parity check matrix can be defined by stacking the stabilizers in the rows of a matrix of the form:
$\mathbf{H}=\begin{bmatrix}s_{1}\\ s_{2}\\ \vdots\\ s_{n-k}\end{bmatrix}.$
The symplectic format allows a binary representation of the above parity check matrix. In this format, each of the $n$ qubit Pauli operators is represented using a $2n$ bit string. For example, one qubit Pauli operator follows;
$X\mapsto(1|0),$ (3)
$Y\mapsto(1|1),$ (4)
$Z\mapsto(0|1).$ (5)
Therefore, the parity check matrix can be described in binary form as:
$\mathbf{H}=\begin{bmatrix}\mathbf{H}_{X}|\mathbf{H}_{Z}\end{bmatrix}.$ (6)
For any valid code, the stabilizers must all commute with each other, requiring the following to hold:
$\mathbf{H}_{X}\mathbf{H}^{T}_{Z}+\mathbf{H}_{Z}\mathbf{H}^{T}_{X}=0.$ (7)
This constrains the development of quantum codes. Particularly in this work, we focus on a very well-studied class of quantum codes called quantum low-density parity check (QLDPC) codes [17]. More specifically, we focus on the class of Calderbank-Shor-Steane (CSS) codes [18, 19] with a special property that the number of $1$’s in each row and column is upper bounded by a constant with the number of qubits going to infinity. This results in a very sparse parity check matrix, which is very efficient with respect to iterative decoding. The stabilizers of these codes are of the independent type, meaning they contain either Pauli-$X$ or Pauli-$Z$ operators, and as a result, Eq. (6) takes a special form
$\mathbf{H}=\begin{bmatrix}\mathbf{H}_{X}&\mathbf{0}\\ \mathbf{0}&\mathbf{H}_{Z}\end{bmatrix}.$ (8)
This ensures that the decoding can be performed independently for syndromes due to bit flips and the syndromes due to phase flips. Also, Eq. (7) can now be reduced to
$\mathbf{H}_{X}\mathbf{H}^{T}_{Z}=0.$ (9)
Following this, we explore a very special class of QLDPC codes, namely the Generalized Hypergraph Product (GHP) codes. It is the generalization of the hypergraph construction, and holds great popularity due to its improved asymptotic performance and better distance scaling.
II-C Generalized Hypergraph Product (GHP) Codes
An intuitive way to construct QLDPC codes based on Eq. (9), is using two commuting matrices and stack them together to form the parity check matrix of the QLDPC code [20]. Panteleev and Kalachev use this intuition to propose a generalization of the standard hypergraph product codes [7]. In their work, they consider two square matrices, elements of which belong to a ring called the ring of circulants defined below in II.1.
Definition II.1 (Ring of Circulants).
A ring of circulant $\mathbb{A}_{L}$ satisfies all the properties of a ring, with its members being the circulant permutation matrices, or sometimes referred to as the polynomials. The whole ring is characterized by a parameter called "lift", which is essentially the dimension of the square permutation matrices.
If the ring of circulants contains only null and identity, then the GHP CSS code reduces to HGP code. The parity check matrix of a GHP code is of the following form:
$\mathbf{H}_{X}=[A,bI_{m}]\,\,\text{and}\,\,\mathbf{H}_{Z}=[b^{T}I_{n},A^{T}],$ (10)
where A is a matrix $A=[a_{ij}]_{m\times n}$, such that $a_{ij}\in\mathbb{A}_{L}$ and $m=n$. Further, $b$ is a polynomial in $\mathbb{A}_{L}$ and $I_{m}$ and $I_{n}$ are diagonal matrices with their diagonal entries being the identity of the ring. For example, we will extensively use the $[[882,24]]$ code discussed in [7], which uses
$A$ $=\begin{pmatrix}x^{27}&0&0&0&0&1&x^{54}\\ x^{54}&x^{27}&0&0&0&0&1\\ 1&x^{54}&x^{27}&0&0&0&0\\ 0&1&x^{54}&x^{27}&0&0&0\\ 0&0&1&x^{54}&x^{27}&0&0\\ 0&0&0&1&x^{54}&x^{27}&0\\ 0&0&0&0&1&x^{54}&x^{27}\end{pmatrix},$ (11)
$b$ $=\begin{pmatrix}1+x+x^{6}\end{pmatrix}I_{7},$ (12)
where $x^{i},\,\,\forall i\in\{1,...,L-1\}$ are the polynomials with it’s columns shifted by $i$ positions, compared to the corresponding identity element. Also, the lift for the ring is set to $L=63$; thus, all these polynomials are $(63\times 63)$ matrices over $\mathbb{F}_{2}$. We also experiment with the $[[1270,28]]$ GHP code (see the appendix of [7] for details).
II-D Iterative Decoding
We discussed in the previous sections that errors anti-commuting with the stabilizers result in a non-zero syndrome, indicating the occurrence of errors. Based on the received syndrome $\mathrm{s}$, the decoding problem is a search for the error, which maximizes the posterior probability:
$\hat{e}=\operatorname*{arg\,max}_{e\in\mathcal{G}_{n}}P(e|\mathrm{s}),$ (13)
where, $\hat{e}$ is the best estimate of the error $e$ and $\mathcal{G}_{n}$ is the set of all $n$-qubit Pauli operators. In case of quantum error correction, the optimal decoding is instead solving the following optimization:
$\hat{e}=\mathrm{any}(E):E=\operatorname*{arg\,max}_{E\in\mathcal{E}}P(E|\mathrm{s}),$ (14)
where $E$ is the coset of an error $e\in\mathcal{G}_{n}$ such that $E:=e\mathcal{S}$. $\mathcal{E}$ is a collection of all such possible cosets. The estimated error $\hat{e}$ of Eq. (14) can be any operator from the maximum probability coset $E$, as members of a single coset have the same logical effect on the state.
BP is an iterative message-passing algorithm due to its continuous back-and-forth information-passing characteristics over a bipartite Tanner graph of the code. BP primarily evaluates the marginals of Eq. (13), i.e.
$\hat{e}_{i}=\operatorname*{arg\,max}_{e_{i}\in\{e_{1},e_{2},....,e_{n}\}/e_{i}}P(e_{i}|\mathrm{s}),$ (15)
Therefore, it so happens that the maximized $\hat{e}$ in (13) does not always satisfy Eq. (14). Many variants of the BP algorithm exist, primarily depending on the application of the algorithm. We describe very briefly the min-sum algorithm, which we used for the numerical experiments, and also offers better numerical stability than the others. This algorithm performs a series of binary message passing from data qubit nodes to its adjacent check nodes and vice-versa, producing soft outputs for each data qubit node. We use $v_{i}$ and $c_{j}$ to denote the $i^{th}$ data and $j^{th}$ check nodes respectively. Also, we assume the notation $m^{k}_{a\rightarrow b}$ to portray the message passed at the $k^{th}$ iteration of the algorithm from node $a$ to node $b$ on the Tanner graph. Following we describe an overall brief summary of the standard message passing rule of the min-sum BP:
  1. 1.
    Initialization: Initially from each data qubits a message $m^{0}_{v_{i}\rightarrow c_{j}}$ is passed as follows:
    $m^{0}_{v_{i}\rightarrow c_{j}}=\lambda_{i}=\log(\frac{1-p_{i}}{p_{i}});\,\,\forall i,j:\mathbf{H}_{ji}=1.$ (16)
  2. 2.
    Check to data qubit messages: For the subsequent iterations, message $m^{k}_{c_{j}\rightarrow v_{i}}$ is passed from each check node to data qubits. Although note we exclude the message passing to the qubit node, which was immediately involved in the same message passing route at ${k-1}^{th}$ iteration. We define a quantity $w=\min_{v_{p}:\,N(c_{j})/v_{i}}\{|m^{k-1}_{v_{p}\rightarrow c_{j}}|\}$ and the message passed is as follows:
    $m^{k}_{c_{j}\rightarrow v_{i}}=(-1)^{\mathrm{s}_{j}}\alpha\left(\prod_{v_{p}:\,N(c_{j})\backslash v_{i}}\mathrm{sign}(m^{k-1}_{v_{p}\rightarrow c_{j}})\right)w,$ (17)
    where, $\alpha$ is a scaling factor and can affect the convergence and $\mathrm{s}_{j}$ is the syndrome value corresponding to the check node $c_{j}$. Also, here and in the rest of the article, we denote the set of neighborhoods of any data qubit $q_{i}$ or stabilizer check $c_{j}$ as $N(q_{i})$, $N(c_{j})$, respectively.
  3. 3.
    Data qubit to check messages: Except for the initialization as described, at each iteration $k$, the messages passed from data qubits to check nodes follows:
    $m^{k}_{v_{i}\rightarrow c_{j}}=\lambda_{i}+\sum_{c_{p}:N(v_{i})\backslash c_{j}}m^{k-1}_{c_{p}\rightarrow v_{i}}$ (18)
  4. 4.
    Hard decision: At the end of each iteration, a soft output is evaluated for each qubit node. The soft output represents the collective information received from all the check nodes and is as follows:
    $\gamma_{i}\leftarrow\lambda_{i}+\sum_{c_{p}:N(v_{i})}m^{k}_{c_{p}\rightarrow v_{i}}.$ (19)
    If $\mathrm{sign}(\gamma_{i})=-1$, then the hard decision for data qubit $i$ is $\hat{e}_{i}=1$, otherwise $\hat{e}_{i}=0$.
The whole process continues until all the unsatisfied checks are satisfied or a pre-determined number of message-passing iterations have been completed, yet the number of unsatisfied checks is still non-zero. In the former, the decoding is said to have succeeded, and the latter reflects a decoding failure. A cycle-free Tanner graph guarantees the success of BP within a finite number of iterations, although QEC offers plenty of short cycles along with inevitable degeneracies, which makes BP fail miserably. In the following section, we discuss such cases where some inherent configurations of the GHP code renders min-sum BP as an ineffective decoder.
$v_{0}$$c_{0}$$c_{1}$$v_{1}$$c_{2}$$c_{7}$$c_{6}$$v_{6}$$c_{12}$
(a) Classical-type Trapping Set
$v_{0}$$c_{0}$$v_{477}$$c_{405}$$v_{405}$$c_{406}$$v_{478}$$c_{352}$$v_{351}$$c_{357}$$v_{483}$$c_{6}$$c_{1}$$c_{411}$$c_{351}$
(b) Quantum Trapping set
Figure 1: In (a) we show a typical $(3,3)$ 6-cycle classical type trapping set and in (b) is a $(6,0)$ Quantum Trapping set of the $[[882,24]]$ GHP code.
III Trapping Sets and Qubit Separation of QLDPC codes
Iterative message-passing-based decoders are vulnerable when it comes to certain error configurations. These are heavily studied and formally named as the Trapping sets (TS). We use $\mathcal{T}_{s}$ to denote the set of data qubits that are trapped in any TS. If $\mathcal{T}_{s}$ contains $a$ data qubits and has $b$ number of odd degree stabilizer checks adjacent to $\mathcal{T}_{s}$, then it is a $(a,b)$ TS of the code. Any $(a,b)$ TS has many possible configurations, depending on the code structure and the decoder under consideration. In the following, we primarily discuss the necessary concepts of TS in the context of GHP codes and show the typical structure of the GHP code trapping sets. Later, we discuss how the notion of qubit separation is exploited to identify the stabilizer check nodes, whose contributions, if ignored, both the trapped qubit’s separation and the iterative decoder’s performance are increased.
III-A Trapping set analysis
For QLDPC codes, there exist two classes of harmful TS configurations: Classical-type trapping sets (CTS) and Quantum trapping sets (QTS). The definitions of CTS and QTS are briefly covered in III.1 and III.2, which we adopt from [12].
Definition III.1 (Classical-type Trapping sets).
A classical-type trapping set is a set of qubits that either do not converge or are adjacent to unsatisfied stabilizer checks after a predetermined number of iterations of a syndrome-based iterative decoder.
For example in Fig. 1(a), we show a $(3,3)$ CTS obtained from the $\mathbf{H}_{Z}$ of the $[[882,24]]$ GHP code discussed in (11). As can be seen from the qubit and check node indexing, the CTS is formed only due to the contributions of a particular circulant matrix of the form $b^{T}I_{n}$. This type of CTS formation is also discussed in [21]. The min-sum iterative decoder in II-D fails to converge in this scenario due to the symmetric message passing rules. For example, for a received syndrome $\mathrm{supp}(\mathrm{s})=\{c_{0},c_{1},c_{2},c_{6},c_{7},c_{12}\}$, indicating the violated stabilizer checks, the min-sum BP based decoder oscillates between an all-zero error pattern and an error pattern of $\text{supp}(e)=\{v_{0},v_{1},v_{6}\}$, as shown in Fig 2.
$v_{0}$$c_{0}$$c_{1}$$v_{1}$$c_{2}$$c_{7}$$c_{6}$$v_{6}$$c_{12}$100010001
(a) Iteration $\mathcal{K}=1$
$v_{0}$$c_{0}$$c_{1}$$v_{1}$$c_{2}$$c_{7}$$c_{6}$$v_{6}$$c_{12}$000000000
(b) Iteration $\mathcal{K}=2$
$v_{0}$$c_{0}$$c_{1}$$v_{1}$$c_{2}$$c_{7}$$c_{6}$$v_{6}$$c_{12}$100010001
(c) Iteration $\mathcal{K}=3$
Figure 2: We show a typical oscillatory behaviour of the min-sum BP algorithm (II-D) for the $(3,3)$ CTS of the $[[882,24]]$ GHP code. The square colour nodes represent the input unsatisfied check nodes. The final hard information passed to each of the data qubits $v_{j}\in V_{1}$ is shown for each complete iteration of the algorithm. Qubits $v_{j}\notin V_{1}$ are not affected by the message passing and thus not highlighted. The algorithm is initialized with errors indicated at $v_{0}$, $v_{1}$ and $v_{6}$. Thereafter, a non-convergent oscillatory behaviour between all zero and $\{v_{0},v_{1},v_{6}\}$ is observed.
The other type of TS that often occurs in the QLDPC codes is the Quantum Trapping sets (QTS) as defined in III.2. These trapping sets do not have any odd-degree check nodes. Thus, a QTS with $a$ number of data qubits is referred as a $(a,0)$ trapping set. Fig. 1(b) shows a typical QTS formed for the same GHP code with $6$ data qubits in the $\mathcal{T}_{s}$.
Definition III.2 (Quantum Trapping sets).
A quantum trapping set is a collection of an even number of qubits and symmetric stabilizers, such that the induced sub-graph contains no odd-degree stabilizer check nodes. This type of trapping set can always be partitioned into isomorphic disjoint subsets with a set of common odd-degree stabilizer (check) neighborhoods.
The two disjoint symmetric qubit subsets of the $(6,0)$ TS, can be identified as $V_{a}=\{v_{0},v_{351},v_{405}\}$ and $V_{b}=\{v_{477},v_{478},v_{483}\}$. Further, as per the definition III.2. $V_{a}$ and $V_{b}$ have the same set of odd-degree stabilizer check neighborhood. That is, $N(V_{a})=N(V_{b})=\{c_{0},c_{1},c_{6},c_{405},c_{406},c_{411},c_{351},c_{352},c_{367}\}$. Following the Lemma 1 from [12], error patterns of cardinality less or equals $3$ on any of the subset is a harmful configuration for the $(6,0)$ TS. For instance any $e_{a}\subseteq V_{a}:|e_{a}|\leq|V_{a}|$ has an exact twin $e_{b}\subseteq V_{b}$. This results into an oscillation between $e_{a}\oplus e_{b}$ and an all zero error at each iterations of the min-sum decoder discussed in II-D.
Lemma 1.
Iterative decoders with critical number $\frac{a}{2}$ posses no TS characteristics for the $(a,0)$ QTS, if the cardinality of the error subsets $e_{a}\subseteq V_{a}$ or $e_{b}\subseteq V_{b}$ exceeds $\frac{a}{2}$.
Note that, the common stabilizer check neighborhood of the $(6,0)$ QTS is itself the set of odd degree check nodes, when we consider the disjoint subsets independently. Therefore, the potential violation of check nodes occur from $N(V_{a})$ or, $N(V_{b})$ (i.e. $\mathtt{UNSAT}\subseteq N(V_{a})=N(V_{b})$) for any of the possible harmful error patterns for the min-sum decoder. This particular phenomena requires different measures compared to the approach for CTS, while attempting to improve the iterative decoder’s performance using qubit separation, which we elaborate next.
III-B Mitigate TS configurations through “qubit separation"
We now show that the iterative decoders can leverage from the increased “separation” of the trapped data qubits. The oscillatory behavior of the TS discussed in the previous section is majorly contributed by the mis-satisfied check nodes of the TS, as discussed in [16]. In the context of CTS for quantum codes, one can assume that the mis-satisfied nodes are the satisfied stabilizer checks, which are adjacent to an even number of incorrect qubits. Therefore, one can relate that these types of nodes are specific to the CTS. In a $(a,b)$ CTS, mis-satisfied check nodes can be present in any amount. For example, in case of the $(3,3)$ TS in Fig 1(a); $c_{1}$, $c_{6}$ and $c_{7}$ are the mis-satisfied check nodes. The term mis-satisfied justifies the dilemma of passing incorrect messages during decoding iterations, although being satisfied. For instance the messages due to $c_{1}$ and $c_{6}$, results in a non-convergence for $v_{0}$. This phenomenon is also reflected in Fig 2. This leads us to a natural intuition, where if we can identify such mis-satisfied stabilizer check nodes, removal of such can increase the possibility of the iterative decoder’s convergence, through a message of correct information.
Kang et al. proposed the use of bit separation for the classical TS [16]. It is essentially a measure of how far an incorrect trapped bit is from another bit trapped in the same TS, found along the computation tree of the former bit [22]. The computation tree (CT) is often used to analyze the performance of iterative decoders in classical settings [22, 23]. We adopt the definition of the CT from [16, 22] and in the context of quantum codes, it is as follows:
Definition III.3 (Computation Tree).
The computation tree $\mathcal{T}_{\mathcal{K}}(t)$ for an iterative decoder on a code is constructed by picking a root node ‘$t$’ corresponding to either a qubit or a stabilizer check node and then iteratively adding ‘$\mathcal{K}$’ levels of edges and leaf nodes according to each complete iterations (i.e., qubit (check) $\rightarrow$ check (qubit) $\rightarrow$ qubit (check)) of the messages passed for the iterative decoder.
We can now explore the role of CT construction in the analysis of qubit separation. Consider a TS $\mathcal{T}_{s}$ having a set of odd-degree data qubits and stabilizer checks $V_{1}$ and $C_{!}$ respectively. We start building the CT from a trapped qubit $v\in V_{1}$ or from an unsatisfied check node $c\in C_{1}$ as a root. From the root, we add descendant stabilizer check or data qubits according to the iterative message passing, which directly follows from the Tanner graph of the code. Due to this, for the rest of the article we use $\mathcal{K}$ to denote the level of CT as well as the number of complete iterations of iterative decoders on similar note. We now define the separation of a qubit trapped inside a CTS:
Definition III.4 (Qubit separation for CTS).
If an erroneous data qubit node $v\in V_{1}$ for any $V_{1}\subset\mathcal{T}_{s}$; has at least one degree-one check node $c$, such that, within $\mathcal{K}$ number of message passing iterations of the decoder, there exists no more data qubit $u\in V_{1}$ as a descendant of $c$ in $T_{\mathcal{K}}(c)$, then the root qubit $v$ is said to be $\mathcal{K}$ separated.
$v_{0}$$c_{0}$$c_{1}$$c_{6}$$v_{477}$$v_{62}$$v_{57}$$v_{513}$$v_{567}$$c_{5}$$c_{62}$$v_{5}$$v_{4}$$c_{6}$$c_{11}$$v_{6}$$v_{10}$$v_{11}$$v_{56}$$c_{57}$$v_{57}$123
Figure 3: Computation tree of the trapped qubit $v_{0}$ of CTS $(3,3)$. We show one of the trapped qubits as a descendant of $v_{0}$ at layer $\mathcal{K}=3$. At each level, the nodes (data or stabilizer check) are added according to where the messages are being passed from parent to child based on the iterative decoder.
In Fig. 3, we show the computation tree $T_{4}(v_{0})$ for the $(3,3)$ CTS of $[[882,24]]$ GHP code. We observe that the qubit $v_{0}$ has a separation of $\mathcal{K}=2$, as at $\mathcal{K}=3$ there exists $v_{6}\in V_{1}$. From Fig. 3, we also observe at level $\mathcal{K}=3$ of $T_{4}(v_{0})$, if we remove $c_{6}$, the separation of $v_{0}$ is increased. We also note that $c_{6}$ is indeed a mis-satisfied node as discussed above. Therefore, an increase in separation directly relates to the removal of mis-satisfied check nodes. Proposition $2$ of [16] or theorem 1 of [24], supports that a sufficiently large separation leads to correctable root trapped qubit of the CT by an iterative decoder. The same can be observed from Table. I.
Removed Stabilizer check nodes Affected trapped qubit New syndrome bit predictions by min-sum decoder
$c_{6}$ $v_{0}$, $v_{6}$ $[c_{0},c_{1},c_{7},c_{12}]$
$c_{1}$ $v_{0}$, $v_{1}$ $[c_{0},c_{2},c_{6},c_{7}]$
$c_{7}$ $v_{1}$, $v_{6}$ $[c_{1},c_{2},c_{6},c_{12}]$
TABLE I: Direct check node removal showing the positive impact on the prediction of erroneous trapped qubits by min-sum BP decoder for a $(3,3)$ CTS.
The nature and harmful effect of QTS is intrinsically different than the CTS. Considering the differences, we propose a slightly different definition for the qubit separation in the context of QTS.
Definition III.5 (Qubit separation for QTS).
If a QTS has degenerate subsets $V_{a}$ and $V_{b}$, then for an erroneous data qubit node $v\in V_{a}$ with atleast one degree-one (w.r.t. the same subset) check node $c$, is said to be $\mathcal{K}$ separated, if within $\mathcal{K}$ number of message passing iterations of the decoder, there exists no more data qubit $u\in V_{a}$ as a descendant of $c$ in $T_{\mathcal{K}}(c)$.
The same definition extends towards the separation of qubits in $V_{b}$. In the context of QTS, the mis-satisfied check nodes do not exist. Instead, a subset of the common stabilizer check neighborhood is responsible for passing incorrect information to the erroneous data qubits in the QTS. These check nodes are present in level $1$ of the computation tree with the erroneous data qubit as it’s root. Removal of these subset of nodes improves the separation of the root erroneous qubit and helps iterative decoder pass correct information to the erroneous qubit.
Lemma 2.
A QLDPC code with data qubits having degree $d_{v}$, has a total of $d_{v}(d_{v}-1)$ number of stabilizer check nodes $c_{r}:c_{r}\in N(V_{a})=N(V_{b})$, which limits the separation of any erroneous data qubit $v_{r}\in V_{a}$ or, $v_{r}\in V_{b}$ with $V_{a}$ and $V_{b}$ together forming the QTS.
Proof.
The proof of the above is very simple and is a direct consequence of the isomorphic property of the QTS. As in a typical QTS, there are no odd-degree stabilizer check nodes; therefore, all the qubits in the QTS exhibit $d_{v}$ number of adjacent stabilizer check nodes. Further from the definition of QTS, we can conjecture that the isomorphic nature of the two disjoint subsets results in a set of qubits at level $2$ of $T(v_{r})\forall v_{r}\in V_{i}$, which limits the qubit separation. We assume these qubits are child nodes of the parent $p$ number of check nodes. Each leaf data qubit nodes at level $1$ of $T(v_{r})$ has $d_{v}-1$ number of child check nodes at level $2$, therefore each of these $d_{v}-1$ number of nodes has one child data qubit in the $T(v_{r})$, which contributes to limiting qubit separation. So, $p=d_{v}-1$. Now finally we need to look at how many parent data qubit nodes at level $1$ of the CT, has these $d_{v}-1$ number of check nodes at level $2$. This number is simply $d_{v}$, because the root qubit has $d_{v}$ number of child check nodes and from each such check nodes one of the descendant qubit in level $1$ belongs to the other isomorphic qubit subset of the QTS. Therefore, in total $d_{v}(d_{v}-1)$ number of check nodes are connected to qubits at level $2$, which contributes to the limiting condition for qubit separation. ∎
Lemma 3.
Any erroneous qubit $v_{r}\in V_{i}$, where $V_{i}$ are the isomorphic qubit subsets of the QTS; can be corrected by min-sum BP decoder if the $d_{v}(d_{v}-1)$ number of check nodes from the level $1$ of the CT, $T(v_{r})$, which are limiting the separation of $v_{r}$; are removed from the code’s Tanner graph.
Proof.
To proof the above lemma, we analyze the message received by one of the erroneous data qubits in the QTS. For instance assume $v_{r}\in V_{i}$, where $V_{i}$ is one of the degenerate qubit subsets of the QTS and $c_{r}\in N(V_{i})$ is one of the common stabilizer check nodes, which connects two different disjoint qubit subsets of the QTS. Also, $T(v_{r})$ is the computation tree with $v_{r}$ as it’s root. We now observe the flow of information in $T(v_{r})$ from leaf nodes towards the root node. For the min-sum BP decoder to be able to correct erroneous $v_{r}$, correct information must be passed from $c_{r}$ to root $v_{r}$ in $T(v_{r})$. For this to be done, all the information passed to $c_{r}$ from it’s descendants must be correct. The correct passage of information must satisfy the following equality for messages passed from any child node $v_{i}$ to the parent $c_{r}$,
$\mathrm{sign}(\lambda_{i})=\mathrm{sign}(m_{v_{i}\rightarrow c_{r}});\forall v_{i}\in N(c_{r})\backslash v_{r}.$ (20)
Eq. (20) imply that as information passed into $c_{r}$ is correct, sign of all the messages passed into $c_{r}$ should be same as that of the message passed into data qubit nodes $v_{i}$ from the channel. Now, assuming phenomenological bit flip noise channel, the magnitude of the messages received by the data qubit nodes from the channel are equal. Therefore for $d_{v}$ regular QLDPC code, the messages send from any data qubit $v_{i}$ in the same level of $T(v_{r})$ has equal magnitude and follows
$|m_{v_{i}\rightarrow c_{j}}|$ $=|\lambda_{i}|+\sum_{c_{j^{\prime}}:N(v_{i})\backslash c_{j}}|m_{c_{j^{\prime}}\rightarrow v_{i}}|.$ (21)
Therefore, if we consider the effect of the stabilizer check nodes at level $2$ of $T(v_{r})$, we directly observe violation of Eq. (20). This is because of the following condition
$\mathrm{sign}(m_{c_{j}\rightarrow v_{i}})\neq\mathrm{sign}(\lambda_{i}),$ (22)
where if $v_{r}\in V_{i}$, then $v_{i}\in V_{j}$; such that $V_{i}\cap V_{j}=\emptyset$. At the junction of level $1$ and level $2$, qubit $v_{i}$ is adjacent to $d_{v}-1$ child check nodes and all such nodes follow Eq. (22). Therefore, the check nodes at level $2$ of $T(v_{r})$, which limit the separation of $v_{r}$, also pass incorrect information to check node $c_{r}$ adjacent to $v_{r}$, which leads to net incorrect information passed to $v_{r}$ by the min-sum BP decoder. Therefore, removal of all such $d_{v}(d_{v}-1)$ number of check nodes from level $2$ of $T(v_{r})$ ensures correct information passed to the erroneous qubits and, as a result, ensures correctness of the erroneous qubits in the QTS. ∎
$v_{0}$$c_{0}$$c_{1}$$c_{6}$$v_{478}$$v_{477}$$v_{483}$$c_{406}$$c_{352}$$v_{405}$$v_{351}$$c_{351}$$c_{405}$$v_{405}$$v_{351}$$c_{411}$$c_{357}$$v_{405}$$v_{351}$12
$v_{405}$$c_{405}$$c_{406}$$c_{411}$$v_{478}$$v_{477}$$v_{483}$$c_{352}$$c_{1}$$c_{357}$$c_{6}$$v_{0}$$v_{351}$$v_{351}$$v_{0}$$c_{0}$$c_{351}$$v_{0}$$v_{351}$12
$v_{351}$$c_{357}$$c_{352}$$c_{351}$$v_{478}$$v_{483}$$v_{477}$$c_{0}$$c_{405}$$v_{0}$$v_{405}$$c_{6}$$c_{411}$$v_{0}$$v_{405}$$c_{1}$$c_{406}$$v_{0}$$v_{405}$12
Figure 4: Computation tree and the potential structure indicating the scope of separation improvement for all the trapped qubits of one of the disjoint subsets of the $(6,0)$ QTS of $[[882,24]]$ GHP code. Note the dotted branches indicated in each of the computation trees. These are the branches bringing flexibility towards separation increments in a QTS enabled by definition III.2. The square check adjacent to leaf blue nodes are the candidate set for removal. The two propositions discussed above can be verified from the diagram.
In Fig. 4, we show the CT with root qubits from one of the degenerate qubit subsets of the $(6,0)$ QTS of the $[[882,24]]$ GHP code. This code belongs to a family of GHP codes with $d_{v}=3$. We validate the results of Lemma 2 and observe that for each of the erroneous trapped qubits, there are exactly $d_{v}(d_{v}-1)=6$ stabilizer check nodes at level $2$ of the computation tree, which are limiting the separation of the trapped qubits.
Removed Stabilizer check nodes Affected trapped qubit New syndrome bit predictions by min-sum decoder
$[c_{406},c_{352},c_{351},c_{405},c_{357},c_{411}]$ $v_{0}$ $[c_{0},c_{1},c_{6}]$
$[c_{0},c_{405},c_{1},c_{406},c_{6},c_{411}]$ $v_{351}$ $[c_{351},c_{352},c_{357}]$
$[c_{0},c_{351},c_{1},c_{352},c_{6},c_{357}]$ $v_{405}$ $[c_{405},c_{406},c_{411}]$
TABLE II: Direct check node removal showing the positive impact on the prediction of erroneous trapped qubits by min-sum BP decoder for a $(6,0)$ QTS.
In table II, we validate the results of lemma 3 and observe that removal of those separation limiting $6$ stabilizer check nodes actively improves the min-sum BP decoder’s performance and ensures correction of the root erroneous trapped qubit.
In the following section, we first discuss an analogy with the classical results from Zhang et al. [24] to discuss the limitations imposed by the probabilistic decoding algorithm. We then use the concepts of “information measurement" from [25] to improve the possibility of an accurate set of stabilizer check removals from the code’s Tanner graph and discuss a new algorithm which provides a better success rate than the min-sum BP algorithm. We perform numerical memory experiments to:
  1. 1.
    show the advantage of removing proper check nodes through a collaborative decoding approach,
  2. 2.
    benchmark the same against min-sum BP and state-of-the-art BP+OSD decoder.
IV Decoding via Collaborative Check Node Removal
As discussed in the previous section, removing the check nodes from the computation tree of an erroneous trapped node increases its separation and improves the performance of min-sum BP decoder. However, knowing the exact level in the CT from where the check node removal can be done requires the TS information, specially for the CTS. Even so, identification of such erroneous trapped qubits also requires the information of TS. For example, Fig. 3 and 4 in Section III-B clearly identify the removable candidate check nodes from the CT of the erroneous trapped qubits. These removable check nodes are present at level $\mathcal{K}=3$ and $\mathcal{K}=2$ of the CT of respective erroneous qubits belonging to the $(3,3)$ CTS and $(6,0)$ QTS, respectively. These candidate check nodes are a set of mis-satisfied check nodes and a subset of common stabilizer check nodes of the CTS and QTS, respectively, information of which can only be available upon knowing about the corresponding TS. Therefore, a natural, practical way to approach this scenario is through the same probabilistic approach of [16], which increases the separation of the trapped nodes without knowing the particulars of the TS of the quantum code. Next, we first discuss the method proposed by Kang et al. in the context of QLDPC codes. We discuss the concept of weak check nodes and how it allows us to derive an algorithm that can improve the min-sum decoder’s performance without knowing the particulars of the TS. Following it, we propose an improved algorithm, namely Quantum Collaborative Check Node Removal (QCCNR); which has more potential to eliminate the vulnerable stabilizer check nodes from the quantum code and achieve better success rates using a collaborative architecture.
IV-A Weak stabilizer check node removal
Here we mainly address a technique relevant to improving min-sum decoder’s performance in the context of CTS. Later, we show that the outline of weak stabilizer check node removal can smoothly merge in the overall decoding algorithm, when we also address the QTS. We saw before that the CTS for the GHP code is rooted in the contributions from a single circulant matrix. For example, the CTS shown in Fig 1(a) is due to the circulant matrix $bI_{7}$. These types of trapping sets have odd-degree check nodes adjacent to them, which we exploit to initialize the decoding. Therefore, the decoding problem is essentially a syndrome-based algorithm, which leverages the CT constructed with unsatisfied check nodes as the roots. We now adopt the notion of week check nodes from [16], which we will show later to be invaluable in the overall decoding architecture. It is defined as:
Definition IV.1 (Weak stabilizer check nodes).
In the computation tree $T_{\mathcal{K}}(c_{r})$ with an unsatisfied check node $c_{r}\in\mathtt{UNSAT}$, as it’s root, has at-least one check node $c$ at any level $t$ of $T_{\mathcal{K}}(c_{r})$, which has the shortest path to any of the nearest qubit $v\in\mathcal{T}_{s}$, then $c$ is a weak stabilizer check node at level $\mathcal{K}=t$ of $T_{\mathcal{K}}(c_{r})$.
Therefore, instead of removing the exact mis-satisfied check nodes from $T(v)$ for any $v\in\mathcal{T}_{s}$, we remove the weak stabilizer check nodes from level $\mathcal{K}=1$ of $T_{\mathcal{K}}(c_{r})\,\forall c_{r}\in\mathtt{UNSAT}$. This improves the separation of the adjacent trapped qubit of the unsatisfied check node $c_{r}$. For example, we observe in Fig 3 that $c_{5}$ is a weak stabilizer check node at $\mathcal{K}=1$ of $T(c_{0})$ for the $(3,3)$ CTS of the $[[882,24]]$ GHP code, shown in Fig. 1(a). This enable us to probabilistically sample the check nodes from level $\mathcal{K}=1$ of $T_{1}(c_{0})$ and a hit on $c_{5}$ increase the separation of the adjacent erroneous qubit $v_{0}$, which in-turn enables the min-sum decoder to predict the same, even if the knowledge of $(3,3)$ CTS is unknown. This introduce a lot of freedom and simplicity to the decoding.
Obviously one can remove weak check nodes from any level of CT $T_{\mathcal{K}}(c_{i})\,\,\forall c_{i}\in\mathtt{UNSAT}$. We chose level $1$ of the CT as the only choice for removing check nodes. This choice corresponds to a universal level in the CT, where removing leaf check nodes serves the interest of both the CTS and QTS. Further, it is obvious that not all the check nodes at any level are the weak nodes and removing these does not have any positive effect on the min-sum decoder. We use the parameter “deselection degree” denoted as $\mathtt{df}$, to sample random check nodes from level $1$ of the CT and later remove them in the process. This method of removing weak leaf nodes randomly is the key assistive part of the overall decoding architecture, which enables the mitigation of harmful configurations without knowing the particulars of the underlying TS. The complete algorithm has a minor modification compared to it’s classical counterpart known as the Fixed Node Check Removal (FNCR) [16]. We call the modified algorithm simply the check node removal (CNR) algorithm, which is described in Algorithm 1.
input : $\mathbf{H}$, $\mathtt{UNSAT}$, $\mathtt{t}$, $\mathtt{df}$
output : Modified parity check matrix $\mathbf{H}_{FN}$
$\mathbf{H}_{FN}=\mathbf{H}$;
  // Initialization
$\mathtt{rem}$ = [];
  // Initialize possible leaf checks to remove
1 for $\mathtt{check}$ in $\mathtt{UNSAT}$ do
2       $\mathtt{leaf}$leftarrow All the leaf nodes of $T_{t}(\mathtt{check})$;
3       $\mathtt{rem}$circ $\mathtt{leaf}$;
[…]
5 foreach $\mathtt{uc}$in $\mathtt{dis}$ do
       $\mathbf{H}_{FN}\leftarrow\mathbf{H}_{FN}\backslash\mathbf{H}[\mathtt{uc}]$;
        // Row (check node) removal
6
return $\mathbf{H_{FN}}$ ;
  // The modified matrix
Algorithm 1 Check Node Removal (CNR)
IV-B Use of ‘Information Measurement’ for better efficacy
In the previous section, we discussed the CNR algorithm inspired by the classical probabilistic FNCR algorithm [16]. As one can observe, CNR acts on a large sample space of possible nodes to probabilistically select candidate check nodes for removal. Here the sample space refers to all the candidates, from which our algorithm randomly choses $\mathtt{df}$ number of stabilizer check nodes for removal. In algorithm 1, this space contains all the elements present in the $\mathcal{K}=1$ level of the CT of an unsatisfied check node. Although this method improves the iterative decoder’s performance within a finite number of iterations [16], for large codes, the number of calls to FNCR or CNR might increase. Further, we observe it can severely harm the success rate of the overall decoder. In this section we attempt to reduce the sample space of check nodes for probabilistic removal. The aim is to exclude any kind of check nodes from the sample space, which are incapable to commit any sort of improvement in the iterative decoder’s performance. This suggests a sample space, elements of which satisfy Lemma 2 regarding any QTS, and for CTS, these elements are the check nodes, which can correct any of the trapped data qubits of the underlying TS. We use the concept of information measurement (IM) to quantify the high value stabilizer check nodes at level $1$ of a CT of a root unsatisfied check node. We claim the nodes with the highest IM values are the check nodes, responsible for limiting the separation of erroneous trapped qubits, especially in the context of QTS. IM associates a metric to both the data qubits and stabilizer checks and is defined as follows:
Definition IV.2 (Information Measurement for data qubits).
For each data qubit of the quantum code, the information measurement is defined by the total number of adjacent stabilizer checks that are unsatisfied.
i.e.
$\mathrm{IM}_{v_{i}}=\sum_{j}c_{j},\forall c_{j}\in N(v_{i}),\text{if}\,\,c_{j}=1.$ (23)
Definition IV.3 (Information Measurement for Stabilizer checks).
For each stabilizer check, the information measurement is defined by the sum of information measurement values of all of its adjacent data qubits.
i.e.
$\mathrm{IM}_{c_{j}}=\sum_{i}\mathrm{IM}_{v_{i}},\forall v_{i}\in N(c_{j}).$ (24)
Consider the CT of $v_{351}$ shown in Fig. 4. Assume a typical harmful error configuration $e$ with support $\mathrm{supp}(e)=\{v_{0},v_{351},v_{405}\}$. Now, if we evaluate the IM values of the stabilizer check nodes at level $2$ of $T(v_{351})$, the $6$ check nodes shown in the figure are the ones with the highest IM values. Therefore, chosing the check nodes with maximum IM values eliminates the possibilities for any of the redundant check node’s presence during the random sampling. In appendix we provide algorithm 4, which describe the function ‘$\mathtt{FIND\_IMs}$’. This function returns the IM values of input stabilizer check nodes stored in a dictionary. We modify the algorithm 1 and propose Quantum Check Node Removal (QCNR) in algorithm 2.
input : $\mathbf{H}$, $\mathtt{UNSAT}$, $\mathtt{t}$, $\mathtt{df}$
output : Modified parity check matrix $\mathbf{H}_{FN}$
$\mathbf{H}_{FN}=\mathbf{H}$;
  // Initialization
$\mathtt{rem}$ = [];
  // Initialize checks removal
1 for $\mathtt{check}$ in $\mathtt{UNSAT}$ do
2       $\mathtt{leaf}$leftarrow All the leaf nodes of $T_{t}(\mathtt{check})$;
3       $\mathtt{leafval}\leftarrow\textnormal{{Find\_IMs}}(\mathtt{leaf})$;
       $\mathtt{maxleaf}=\mathrm{MAX}_{c}(\mathtt{leafval})$;
        // Check node with maximum IM value
4       $\mathtt{rem}$circ $\mathtt{maxleaf}$;
[…]
6 foreach $\mathtt{uc}$in $\mathtt{dis}$ do
       $\mathbf{H}_{FN}\leftarrow\mathbf{H}_{FN}\backslash\mathbf{H}[\mathtt{uc}]$;
        // check node removal
7
return $\mathbf{H_{FN}}$ ;
  // The modified matrix
Algorithm 2 Quantum Check Node Removal (QCNR)
This algorithm proceeds with selecting the keys with maximum value from the output of the algorithm ‘$\mathtt{FIND\_IMs}$’. These keys are the stabilizer check node indices with the highest IM values compared to all the leaf nodes of the CT. These selected check nodes then form the sample space of nodes, from which sampling of removable stabilizer check nodes is done according to the specified deselection degree, and ultimately, the quantum code’s Tanner graph is modified. Next, we describe the complete collaborative decoding architecture.
IV-C Collaborative architecture
We now leverage both the concepts of ‘weak stabilizer check nodes’ and ‘information measurement’ to rescue the min-sum decoder from the trapped scenarios. In previous sections, we discussed the improvement in the separation of trapped data qubits due to the modified Tanner graph obtained from CNR, in Algorithm 1. Then, we introduced QCNR, in Algorithm 2, which has an even better probability of having trapped qubits with increased separation, therefore offering better convergence with iterative decoders.
We now propose the use of QCNR in a collaborative setting to deduce a low complexity decoding architecture. We mainly use the iterative or, min-sum decoder in two modes. The main mode implements the min-sum BP on the unmodified parity check matrix of the QLDPC code. The other mode, which we call sub-decoding mode, first removes some stabilizer check nodes from the quantum code’s Tanner graph using Algorithm 2. The min-sum BP is then implemented on the resultant modified parity check matrix of the QLDPC code. The decoding process is initialized with the main decoding mode. As we discussed extensively in previous sections, the main decoding mode is limited in its performance. Therefore, if for a predetermined number of iterations of BP, no more correct predictions are done by the decoder, it switches to the sub-decoding mode. After the main and sub decoding rounds are completed, the decoder switches again to the main decoding mode. We note that in any round of the main and sub decoding rounds, we consider the corrections of all previous rounds, i.e. we continuously update the list of the net unsatisfied stabilizer check nodes. Therefore, calling the main decoding mode again after the completion of a sub-decoding mode is to correct any error that does not require any further sub-decoding (this is meant to tackle errors that are no longer supported over any TS after the previous main+sub-decoding round). This approach is different than the post-processing based methods, as it switches its message-passing operation over different parity check matrices and outputs a prediction only if all the errors are corrected or a predetermined number of sub-decoding rounds have been completed. We describe the complete details of this collaborative decoding in Algorithm 3.
input :
  • Syndrome:= $\mathtt{s}$ and corresponding $\mathbf{H}$,
  • Channel llr values:= $\mathtt{llr}$,
  • Maximum iterations of main BP:= $\mathtt{max_{iter}}$,
  • Maximum iterations of sub BP:= $\mathtt{max_{sub}}$,
  • Deselection degree: $\mathtt{df}$,
  • Maximum sub-decoding rounds: $\mathtt{fr}$,
  • Tolerance of main BP iterations: $\mathtt{tol}$.
output : Predicted error pattern $\mathtt{\hat{e}}$.
1 $\mathtt{unchanged=0}$;
$\mathtt{itn}=1$;
  // Main decoder initialization
2 $\mathtt{s_{prev}:=s}$;
[…]
4       $\mathtt{\hat{e}_{bp}}$ = $\mathsf{BP}$($\mathbf{H},\mathtt{llr},\mathtt{itn}$).$\mathsf{decode}$($\mathtt{s}$);
5       $\mathtt{\hat{s}_{bp}}$ = $\mathtt{\hat{e}_{bp}}\mathbf{H}^{T}$;
6       $\mathtt{itn}\leftarrow\mathtt{itn}+1$;
7       if $\mathrm{supp}(\mathtt{\hat{s}_{bp}})=\mathrm{sup}(\mathtt{s_{prev}})$ then
8            $\mathtt{unchanged\leftarrow unchanged+1}$
9      else
10            $\mathtt{unchanged=0}$
11      $\mathtt{s_{prev}\leftarrow\hat{s}_{bp}}$;
12
$\mathtt{e_{net}}:=\mathsf{zeros}\mathtt{((}\mathtt{fr},\mathbf{H}.\mathsf{shape}\mathtt{[0]}\mathtt{))}$;
  // Sub-decoder init
13 $\mathtt{s_{net}}:=\mathsf{zeros}\mathtt{((}\mathtt{fr},\mathbf{H}.\mathsf{shape}\mathtt{[0]}\mathtt{))}$;
14 while $\mathtt{i<fr}$ do
15       $\mathtt{UNSAT=[]}$;
16       $\mathtt{s_{fn}\leftarrow s+\hat{s}_{bp}+(\sum_{j}s_{sub}[j]),\forall j<i}$;
17       $\mathtt{UNSAT}$.$\mathsf{extend}$($\mathrm{supp}(s_{fn})$);
18
19      $\mathbf{H}_{FN}=\mathsf{QCCNR}(\mathbf{H},\mathtt{UNSAT},\mathtt{t},\mathtt{df})$;
20       $\mathsf{BP_{sub}}=\mathsf{BP}(\mathbf{H}_{FN},\mathtt{llr},\mathtt{max_{sub}})$;
21       $\mathtt{\hat{e}_{sub}=}\mathsf{BP_{sub}.decode}\mathtt{(s_{fn})}$;
22       $\mathtt{s_{sub}\leftarrow\hat{e}_{sub}}\mathbf{H}^{T}$;
23       $\mathtt{\hat{e}^{\prime}_{bp}}$ = $\mathsf{BP}$($\mathbf{H},\mathtt{llr},\mathtt{max_{iter}}$).$\mathsf{decode}$($\mathtt{s_{sub}+s_{fn}}$);
24       $\mathtt{\hat{s}^{\prime}_{bp}}$ = $\mathtt{\hat{e}^{\prime}_{bp}}\mathbf{H}^{T}$;
25       $\mathtt{e_{net}}[i]\leftarrow\mathtt{e_{sub}}+\mathtt{\hat{e}^{\prime}_{bp}}$;
26       $\mathtt{s_{net}}[i]\leftarrow\mathtt{s_{sub}}+\mathtt{\hat{s}^{\prime}_{bp}}$;
27       if $\mathrm{supp}(\mathtt{s_{net}}[i])=\emptyset$ then
28             return $\mathtt{\hat{e}\leftarrow\hat{e}_{bp}+\sum_{m:m\leq i}\mathtt{e_{net}}[m]}$;
29             End;
30
31      $\mathtt{i}\leftarrow\mathtt{i}+\mathtt{1}$
return $\mathtt{\hat{e}\leftarrow\hat{e}_{bp}+\sum_{i}\hat{e}_{net}[i]}$
Algorithm 3 Quantum Collaborative Check Node Removal (QCCNR) Algorithm
In Fig. 5, we show the advantage of using BP in the sub-decoding mode. We observe, out of numerous sampled instances, a case where the main BP decoder gets stuck while decoding the syndrome of an error pattern generated at an independent bit flip physical error rate of $p=0.03$ for the $[[882,24]]$ GHP code after a few iterations. We note that for all the numerical experiments of GHP codes, we use the following parameter values: maximum-BP-iterations = $100$, BP-method = “minimum-sum”, scaling-factor = $0.625$. These parameters are relevant to the message passing of the min-sum BP in both the main and sub-decoding modes. The stuck scenario represents the fact that the main decoder cannot identify any more correct stabilizer check violations. After the main mode of BP decoder fails to correct any more errors for a consecutive $11$ rounds, we iteratively apply Algorithm 2 and run the BP decoder on a modified parity check matrix. We keep track of the new corrections at each round of the sub-decoding. The tail of the plot indicates that by activating the sub-decoding mode, new stabilizer violations are predicted correctly and thus, eventually, all the errors are corrected. This numerically confirms that removing stabilizer check nodes using Algorithm 2 increases the efficacy of the decoding output.
Figure 5: Numerical experiment showing the sub-decoder’s ability to identify stabilizer violations correctly, which were not possible by decoding with min-sum based BP alone (which we denote as ‘main decoding region’). We assume that the error occurs on the data qubits after each round of an error correction cycle, and the syndrome measurements are perfect. Under these phenomenological noise assumptions, we sample an error pattern at an independent bit flip physical error rate of $p=0.03$ for the $[[882,24]]$ GHP code. The min-sum BP decoder can not predict any more correct stabilizer violations for a consecutive $11$ iterations and gets stuck at the $40^{th}$ decoding iteration. Subsequently, we initialize the sub-decoding mode. We call Algorithm 2 repeatedly, and min-sum BP is applied on the output parity check matrix of the QCNR algorithm. To tackle all the trapped cases for QTS and CTS, we first set the deselection degree $\mathtt{df}=6$ and then set $\mathtt{df}=1$ for the last 20 sub-decoding iterations. We observe repeated calls of Algorithm 2 lead to an improvement in the performance of BP in the sub-decoding region. The tail in the plot signifies newly accurate predictions of the violated stabilizer checks and therefore indicates the improved performance.
Further, we perform memory experiments to compare the performance of the proposed QCCNR decoder in Algorithm 3 with the min-sum BP decoder and the state-of-the-art BP+OSD decoder with the order of OSD set to 00. For the decoding of GHP codes in the memory experiments, we used a strategy where we set the maximum number of sub-decoding rounds to $200$. This implicitly includes the case where, if there are no more errors to correct, the decoder stops and outputs the prediction, which is also indicated in Algorithm 3. Now, we have adopted a scenario where in the first $100$ sub-decoding rounds we set the deselection degree to $\mathtt{df}=6$ and for the rest of $100$ rounds we set $\mathtt{df}=1$. This strategy is adopted deliberately to first target the errors supported on any QTS and then to target the errors supported on the CTS. The degree of the data qubits is $d_{v}=3$, for the family of GHP codes we chose. Therefore, from Lemma 2, it is clear that to have the best possible result, we need to set the deselection degree to $\mathtt{df}=d_{v}(d_{v}-1)=6$. In Fig 6, we show the results, which indicate that the proposed decoding scheme offers significant improvements in the logical error rates over the min-sum BP decoder, indicating a potential breakthrough to circumvent the harmful configurations for the iterative min-sum BP. Algorithm 3 gives logical error rates much better than standard min-sum BP and almost achieves a performance comparable to the OSD00 post-processor.
Figure 6: The logical error rates obtained for the GHP codes from [7], under the independent Pauli-$\mathrm{X}$ noise channel. We perform the code capacity LER simulations for both cases. In (a) we benchmark with the $[[882,24]]$ GHP code and in (b) we benchmark the $[[1270,28]]$ GHP code. In both the experiments we use the min-sum algorithm of BP and for benchmarking against the OSD decoder, we use the the $0^{\text{th}}$ order OSD, i.e. BP+OSD00. For the QCCNR algorithm, we use the following parameters for Algorithm 3: $\mathtt{max_{iter}}=\mathtt{max_{sub}}=100,\mathtt{fr}=200$. Also, for the first $100$ sub-decoding rounds, the deselection degree is set as $\mathtt{df}=6$. This is set from Lemma 3 to support qubits trapped inside QTS. Further, for the remaining $100$ sub-decoding rounds, the deselection degree is set to $\mathtt{df}=1$, in support of the qubits trapped inside CTS, which follows from the section IV-A. We observe that the proposed QCCNR algorithm provides OSD-like improvements and obviously is far superior to the standard min-sum algorithm-based BP decoder.
IV-D Time Complexity
Algorithm 3 has two modes, as discussed previously. In both main and sub-decoding mode, the message passing of the BP decoder is linear, i.e., a $\mathcal{O}(n)$ process. In sub-decoding mode, Algorithm 2 has two parts: one is to construct the computation tree for each of the unsatisfied check nodes, and the other is to calculate the IM values and store them in a dictionary. The IM values evaluated for a particular node are an $\mathcal{O}(n)$ process and can be stored once. Essentially, for each round of the sub-decoding, we only require the IM values for a constant number of leaf check nodes, in the CT of each of the remaining unsatisfied check nodes. Therefore, the whole IM value evaluation takes at most $\mathcal{O}(|\mathtt{UNSAT}|n)$ amount of time. For GHP codes this time amounts to $\mathcal{O}(\frac{n^{2}}{\log{n}})$ [2]. The other process involved in Algorithm 2 is the construction of a level $1$ computation tree for each of the unsatisfied check nodes. If the QLDPC code is $(d_{v},d_{c})$ regular, constructing a level $1$ computation tree from any particular root check node requires first adding the $d_{c}$ number of data qubits and then for each of those data qubits adding $d_{v}$ number of stabilizer check nodes. This whole process is a $\mathcal{O}(n)$ process, as $d_{v},d_{c}<<n$ and are constants. Therefore, the overall cost of the decoder is dependent on the maximum number of unsatisfied stabilizers as indicated by the syndrome. For GHP codes, we approximate that this will result in a sub-quadratic decoder. In any case, if $|\mathtt{UNSAT}|$ grows linearly, then the decoder is quadratic. Therefore, QCCNR provides an excellent strategy for improving the iterative decoder’s performance compared to decoders like BP+OSD0 or the higher order alternatives.
V Conclusion and Future Work
In this work, we propose a collaborative decoding approach for improving the iterative decoding of QLDPC codes through improving qubit separation, with a special focus on GHP codes. We propose a new way to measure the separations of the qubits trapped inside a symmetric stabilizer set. We used this method to go beyond the analysis of qubit separation of the classical trapping sets. Our decoding architecture is free of any post-processing method and can be considered as a two-mode decoder with its operation switching the message passing operation of BP over the complete and a modified Tanner graph (or parity check matrix) of the quantum code. The overall decoder’s complexity depends on the support of the syndrome (an implicit dependence on the physical error rate and properties of the code). For GHP codes, it approximately renders a sub-quadratic cost, which, in light of the performance-cost tradeoff, emerges as a practical alternative to the OSD-based post-processing decoders. We showed that, despite being a post-processing-free decoding, the success of the algorithm has a significant improvement compared to the standard iterative min-sum decoder. We would like to conjecture that the QCCNR algorithm does not address the cases of point-like defects. The improved qubit separation can assist only in breaking the trapping set scenarios and is incapable of assisting the iterative decoders when point-like syndrome occurs. Therefore, the slight shift in the decoding performance between QCCNR and BP+OSD00 is due to such point-like syndromes, which might be more reflected in experiments involving small codes like the surface codes. In future work, we will investigate the possible improvements to deal with such point-like defect scenarios and further improve the collaborative architecture’s performance.
Data Availability
The data and scripts used for the numerical experiments are available from the authors on reasonable requests.
Acknowledgements
MB acknowledges ldpc-repository by Joschka Roffe [26], for sourcing BP and BP-OSD [27] decoders used in the memory experiments. AR and MB thank IISER Bhopal for providing the computational resources to conduct the exhaustive simulations. MB acknowledges the doctoral research fellowship from IISER Bhopal.